### v1

# Covariate Adjusted Response Adaptive Design with Delayed Outcomes

Abstract
Covariate-adjusted response adaptive (CARA) designs have gained widespread adoption for their clear benefits in enhancing experimental efficiency and participant welfare. These designs dynamically adjust treatment allocations during interim analyses based on participant responses and covariates collected during the experiment. However, delayed responses can significantly compromise the effectiveness of CARA designs, as they hinder timely adjustments to treatment assignments when certain participant outcomes are not immediately observed. In this manuscript, we propose a fully forward-looking CARA design that dynamically updates treatment assignments throughout the experiment as response delay mechanisms are progressively estimated. Our design strategy is informed by novel semiparametric efficiency calculations that explicitly account for outcome delays in a multi-stage adaptive experiment. Through both theoretical investigations and simulation studies, we demonstrate that our proposed design offers a robust solution for handling delayed outcomes in CARA designs, yielding significant improvements in both statistical power and participant welfare.
Keywords: Causal inference; Delayed outcomes; Frequentist adaptive experimental design; Response adaptive designs
1 Introduction
1.1 Motivation and our contribution
In clinical trials, field experiments in social sciences, and A/B tests, participants often enroll sequentially, and responses to treatments can vary due to individual-specific characteristics. In these contexts, covariate-adjusted response-adaptive (CARA) designs have gained increasing attention for their ability to enhance both the efficiency and the ethical or welfare considerations of experimental outcomes (Zhang et al., 2007; Hu et al., 2015; Robertson et al., 2023; Wei et al., 2023, 2024; Zhao, 2023). By dynamically adjusting treatment assignments based on accumulated responses and covariates during the experiment, CARA designs can be adopted to optimize treatment allocation to improve statistical power, reduce participant exposure to less effective treatments, enhance overall welfare, and maintain statistical validity for analyzing experimental data.
Nevertheless, because CARA designs rely on participant responses collected during the experiment to adjust treatment assignments, delays in observing these responses can compromise the design’s effectiveness. This can be intuitively understood as follows: during interim analyses, if the primary outcomes for some participants are not yet observed, the experimental designer may lack the necessary information to optimize treatment assignments for future participants who have profiles similar to those with delayed responses. Furthermore, as is often the case in randomized experiments when participants are sequentially enrolled, the response delay mechanisms can vary among participants depending on their assigned treatment arm and baseline covariates. Below, we provide a motivating example with arm- and covariate-dependent response delays in a randomized experiment:
Example 1 (Arm- and covariate-dependent randomized experiments for a cash incentive trial among people living with HIV (PLHIV)).
In a randomized experiment conducted in Shinyanga region, Tanzania (Fahey et al., 2020), the goal is to assess the impact of cash incentives on improving retention in care and thus achieving viral suppression for PLHIV. In this experiment, 530 PLHIV aged 18 years or older were sequentially enrolled and randomly assigned to either receive cash incentives (treatment arm) or no cash incentives (control arm) between April 24, 2018, and December 14, 2018. Retention in care is evaluated by individual adherence to antiretroviral therapy measured by viral load six months after treatment assignment. However, delays in measuring viral load are observed, as the lab test requires clinical visits for blood draws, and some PLHIV miss their scheduled appointments. To investigate these response delay mechanisms, we record the date at which each patient completed six months of treatment and calculate the number of days delayed as the difference between the date of the viral load test and the six-month completion date. We further include biological sex as a baseline covariate to examine if the delay mechanism varies across females and males.
Figure 1: Distribution of the number of delayed days in viral load response collection across treatment arms and baseline covariate (i.e., biological sex) after completing six months of antiretroviral therapy.
Figure 1 presents a histogram showing the number of delayed days for viral load measurements, stratified by treatment and control arms as well as biological sex. The figure illustrates that the distribution of delayed days is influenced by both treatment assignment and biological sex. Specifically, the treatment arm shows a higher percentage of responses collected within 100 days compared to the control arm. Moreover, within the treatment arm, a larger proportion of female participants have response delays of fewer than 100 days compared to males. This observation suggests that for females receiving cash transfers, most viral load responses were collected within three months of completing the treatment. Overall, Figure 1 highlights that response delays in this randomized experiment are dependent on both the treatment arm and covariate.
To our knowledge, although there has been discussion about how response delays might affect statistical inference at the conclusion of CARA(Bai et al., 2008; Hu et al., 2008; Robertson et al., 2023), no previous CARA design has leveraged the delay mechanism to explicitly optimize experimental objectives. In this manuscript, we address the challenge of response delays in CARA and seek to answer the following questions: Does the optimal treatment allocation proposed in existing CARA literature remain effective for achieving design objectives? If not, what is the optimal treatment allocation? How can CARA be adapted to incorporate response delays for more effective attainment of these objectives? Finally, at the conclusion of the experiment, how can we construct an efficient statistical estimator that accounts for these delays? The following paragraphs outline our contributions in addressing these questions within the context of existing CARA literature.
First, we demonstrate that the classical optimal allocation strategy used in existing CARA designs no longer provides optimal treatment assignment when accounting for arm- and covariate-dependent delays. Specifically, we first derive a semiparametric efficiency bound for treatment effect estimation accommodating both multi-valued treatments and outcome delay, which differ substantially from existing results in the literature (Hahn, 1998; Hirano et al., 2003; Cattaneo, 2010; Zhu and Zhu, 2023) (see Newey 1990; Van der Vaart 2000; Tsiatis 2006 for general discussions on semiparametric efficiency). By deriving semiparametric efficiency bounds that incorporate response delays in multi-stage CARA designs (Theorem 1), we show that the optimal delay mechanism involves a unique trajectory of treatment assignment probabilities that vary across different stages, resulting in treatment allocations that reduce variance for power maximization and minimize failure rates in welfare improvement designs (Figure 2). To our knowledge, this is the first design to explicitly explore delay mechanisms to optimize trial objectives, while existing literature focuses on conducting sequential tests (Hampson and Jennison, 2013), designing Bayesian adaptive designs where delay is independent of the arm and covariates (Lin et al., 2020), or conducting statistical inference after experiments with delayed responses have concluded (Bai et al., 2008; Hu et al., 2008).
Second, because the delay mechanism is unknown a priori and only partially learnable during the experiment, we propose a fully forward-looking CARA design that sequentially updates treatment assignment probabilities to optimize the trial outcome by the experiment’s end while also estimating future delays to incorporate the experimental designer’s preferences. Our design encounters two design objectives frequently adopted in current CARA literature: maximizing experiment power (Tymofyeyev et al., 2007; Zhao, 2023) and improving overall participant welfare (or failure reduction) while maintaining statistical power constraints (Rosenberger et al., 2001; Robertson et al., 2023). We justify the benefits of our proposed forward-looking CARA design through theoretical investigation (Section 4) and a synthetic case study (Section 5). Under the design objective of maximizing experiment power, our proposed forward-looking CARA design attains higher estimation efficiency compared to the Neyman allocation and complete randomization designs in the presence of delayed responses. The efficiency gain is further justified by Theorem 5. Under the design objective of improving overall participant welfare, we demonstrate that our proposed forward-looking CARA design achieves a higher failure reduction rate, leading to greater welfare improvement compared to both the ethical design and the complete randomization design while maintaining statistical power. For both design objectives, our proposed design can provide valid statistical inference, which is justified in Theorem 2 and Theorem 3 and further demonstrated in a synthetic case study (Section 5).
Third, we make theoretical contributions to the literature on CARA and adaptive experimental designs by establishing general conditions under which estimated treatment effects are consistent and asymptotically normally distributed (Lemma 1 and Theorem 2). Specifically, we do not restrict the potential outcomes or the delay mechanism to any parametric family of distributions, thereby alleviating the burden of choosing a particular set of parametric assumptions. Such flexibility is crucial for many applications, as the delay mechanism is typically arm- and covariate-dependent, as our earlier examples have demonstrated. Our theoretical development builds on martingale methods (Hall and Heyde, 1980) and a high-level condition (Condition 1) that the optimized allocations converge in large samples—an assumption satisfied by many adaptive experimental designs. As a result, the general consistency and asymptotic normality findings may be of independent interest. We then specialize in our proposed design and verify the design consistency condition (Section 4.3), providing a suite of tools for designing and analyzing adaptive experiments with delayed outcomes, along with rigorous statistical guarantees.
1.2 Setup: CARA with arm and covariate dependent delayed outcomes
In the remainder of this introduction, we lay out the setup of CARA when responses are subject to delay. Formal statements of our assumptions are collected in Section 4. The experiment is conducted over $T$ stages, labeled $t=1,2,\dots,T$. In each stage $t$, $n_{t}$ participants are enrolled. Participants are indexed by $i=1,2,\dots,N$, with $N=\sum_{t=1}^{T}n_{t}$ being the total sample size. Occasionally, we will use $N_{t}$ to denote the cumulative sample size at the end of stage $t$, that is, $N_{t}=\sum_{s=1}^{t}n_{s}$. Therefore, $N=N_{T}$, where we omit the subscript to save notation. Finally, $S_{i}=t$ denotes that individual $i$ is enrolled in stage $t$.
Upon enrollment, baseline covariates $X_{i}\in\mathcal{X}$ are collected for each participant, and they are then randomized into one of the treatment arms, denoted by $A_{i}\in\mathcal{A}$, which represents their actual treatment status. Following the Neyman-Rubin causal model, the potential outcomes are denoted by $Y_{i}(a)$ for different treatment $a\in\mathcal{A}$. In line with existing literature in adaptive design, we are in a scenario that the potential outcomes and the covariates, $(Y_{i}(a),X_{i})$, are independently and identically distributed across different stages. For future reference, we also introduce notation for mean potential outcomes: $\mu(x,a)=\mathbb{E}[Y_{i}(a)|X_{i}=x]$ and $\mu(a)=\mathbb{E}[Y_{i}(a)]$, [^1] Other moments of the potential outcomes are $\mu_{s}(x,a)=\mathbb{E}[Y_{i}(a)^{s}|X_{i}=x]$ and $\mu_{s}(a)=\mathbb{E}[Y_{i}(a)^{s}]$, but we omit the subscript whenever $s=1$ for ease of notation. and the conditional and unconditional treatment effects are defined as
$\tau(x,a^{\prime},a)$ $=\mu(x,a^{\prime})-\mu(x,a)\quad\text{and}\quad\tau(a^{\prime},a)=\mu(a^{\prime})-\mu(a).$
Since outcome information may not be immediately available after treatment assignment, we use $D_{i}$ to represent the number of stages after which the experimenter can observe $Y_{i}$. This means that the participant outcome $Y_{i}$ is observed at the end of stage $t$ if and only if $D_{i}+S_{i}\leq t$. To give a concrete example, Let $S_{i}=1$ so the participant is enrolled in the first stage. Then, their outcome information is available at the end of Stage 3 if and only if the delay is no more than two stages, which means $D_{i}\leq 2$, or equivalently, $S_{i}+D_{i}\leq 3$. Also note that when $D_{i}=0$, it implies no delay in the outcome for this individual.
In this manuscript, we allow the delay mechanism $D_{i}$ to depend on the covariates and the treatment assignment status, and is independent of the potential outcomes conditional on the covariates $X_{i}$ and the treatment assignment $A_{i}$. Similar to the assumption of identical distribution of potential outcomes across stages mentioned earlier, we assume that the delay distribution does not shift over time. Together, these can be formally expressed as:
[…]
As the delay distribution does not shift over time, it allows us to sequentially learn the delay mechanism using accrued data and to optimize treatment allocation. Finally, we note that this assumption does not preclude the possibility that the delay mechanism may depend on the observed outcomes.
In CARA designs, the treatment assignment is sequentially updated based on accumulated data to achieve a pre-specified design objective. Specifically, at stage $t$, suppose we have collected historical participant information, denoted as
$\mathcal{H}_{t}=\{(A_{i},X_{i},Y_{i},D_{i})\}_{i=1}^{N_{t}},$
where we recall that the sample indexed by $i=1,2,\dots,N_{t}$ consists of all participants enrolled in the first $t$ stages. The treatment assignment for the next stage, $t+1$, denoted as $\hat{e}^{*}_{t+1}(a|x)$, is determined by both this historical information and the covariate information of the newly enrolled participants in Stage $t+1$. Formally, this is captured by
$\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t+1,\mathcal{H}_{t}]=\hat{e}^{*}_{t+1}(a|x),$
which means the treatment assignment adapts to $\mathcal{H}_{t}$.
2 Optimal delay-adjusted treatment allocation in CARA
In this section, we discuss how delayed response in CARA affects optimal treatment allocation rules. Compared to the existing literature on CARA or response adaptive design (RAR) with delayed outcomes (Bai et al., 2008; Hu et al., 2008), to our knowledge, no prior designs have explicitly leveraged delay mechanisms to optimize experimental objectives.
While the design goals of CARA vary across applications, maintaining statistical power to assess the treatment’s effectiveness on the primary outcome remains a central concern. A statistically efficient estimator of the treatment effect also contributes to a robust adaptive design, as the estimated treatment effects are used to sequentially revise the treatment allocation strategy. However, it remains unclear what constitutes a good estimator of the treatment effect in the presence of delayed outcomes. Moreover, the semiparametric efficiency bound for estimating the treatment effect in multi-stage experiments with delayed outcomes has not been developed. To provide practical guidance for designing response-adaptive experiments, we first present a new result on the semiparametric efficiency bound of estimating the average treatment effect, defined as the mean difference between the potential outcomes of two treatment arms $a$ and $a^{\prime}$. In such experiments with delayed outcomes, we have
$\int\limits_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,a^{\prime})}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a^{\prime}|x)\rho(T-t|x,a^{\prime})}$ $+\frac{\sigma^{2}(x,a)}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}+\big{(}\tau(x,a^{\prime},a)-\tau(a^{\prime},a)\big{)}^{2}\Bigg{]}\mbox{d}x,$ (1)
where $r_{t}=\mathbb{P}[S_{i}=t]$, the fraction of participants enrolled in stage $t$, $p(x)$ is the probability density function of the covariate $X$, and $\sigma_{t}^{2}(x,a)$ is the conditional variance of potential outcomes under treatment $a\in\{0,1\}$ at time $t$. See Section 4, Theorem 1 for a rigorous statement.
From the above result, it is evident that when the delay mechanism of the response depends on the covariates and treatment status, the optimal treatment allocation rules in classical CARA designs must be modified to account for the impact of the delay mechanism. In what follows, we examine how the optimal treatment allocation rule is affected by the arm- and/or covariate-dependent delayed responses in two common experimental objectives: one aimed at enhancing statistical power (Tymofyeyev et al., 2007), and the other focused on minimizing the expected number of failures while maintaining a minimum power requirement for binary outcomes (Rosenberger et al., 2001). In both design goals, the variance of the estimated treatment effects plays an essential role. To simplify the presentation, we focus on two-arm CARA designs with $a\in\{0,1\}$. For multi-arm experiments, the design objective can be generalized using noncentrality parameters, as discussed in Lachin (1977); Tymofyeyev et al. (2007).
Design objective 1: Optimal treatment allocation in power maximization design with delayed response. This design aims to sequentially assign treatments to minimize the variance of the estimated average treatment effect (ATE) for two treatment arms, denoted by $a\in\{0,1\}$, thereby maximizing statistical power. The optimal allocation is determined by solving the following optimization problem:
$\min_{e_{1},\ldots,e_{T}}\ \mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))\quad\text{subject to}\quad\delta\leq e_{t}(\cdot)\leq 1-\delta\quad\text{for }t=1,\ldots,T,$ (2)
where $\delta$ ensures that the assignment probabilities are bounded away from 0 and 1, maintaining sufficient randomness in treatment assignments. The objective function $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ is a component of the semiparametric efficiency bound related to the propensity scores, and is defined as:
$\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))=\sum_{x\in\mathcal{X}}p(x)\left[\frac{\sigma_{t}^{2}(x,1)}{\sum_{t=1}^{T}r_{t}e_{t}(1|x)\rho(T-t|x,1)}+\frac{\sigma_{t}^{2}(x,0)}{\sum_{t=1}^{T}r_{t}e_{t}(0|x)\rho(T-t|x,0)}\right].$
In line with existing literature on CARA designs, we explicitly consider discrete covariates in the objective function, as most design applications will first group participants based on their covariate information. Therefore, the objective function takes a summation form with respect to the probability mass function.
Design objective 2. Optimal treatment allocation in design on failure reduction under a power constraint. This design aims to minimize the expected number of failures while satisfying a minimum power requirement for binary responses (Rosenberger et al., 2001). In this context, a treatment is considered to fail when the potential outcome equals one. The optimal allocation thus seeks to assign treatments based on the solution to the following optimization problem:
$\min_{e_{1},\ldots,e_{T}}$ $\ \sum_{t=1}^{T}r_{t}\sum_{x\in\mathcal{X}}p(x)\Big{[}e_{t}(1|x)\left(1-\mu(x,1)\right)+\left(1-e_{t}(0|x)\right)\left(1-\mu(x,0)\right)\Big{]}$ (3)
s.t. $\ \delta\leq e_{t}(\cdot)\leq 1-\delta,\quad\text{for }t=1,\ldots,T$
$\ \mathsf{V}\left(e_{1}(\cdot),\dots,e_{T}(\cdot)\right)+\sum_{x\in\mathcal{X}}p(x)\big{(}\tau(x,1,0)-\tau(1,0)\big{)}^{2}\leq C$
In this formulation, the objective function minimizes the total expected number of failures during the entire experiment. The term inside the summation accounts for the expected failures when assigning treatment 1 with probability $e_{t}(\cdot)$ and treatment 00 with probability $1-e_{t}(\cdot)$. Here, $\mu(x,a)$ represent the expected success probabilities under treatments $a$, given covariates $x$. The first feasibility constraints again ensure that the treatment assignment probabilities remain within the interval $[\delta,1-\delta]$. This design balances the ethical imperative to reduce adverse outcomes with statistical consideration to maintain sufficient power for detecting treatment effects.
Due to the delay mechanism’s influence on the variance lower bound, the optimal treatment allocation rules differ substantially from those derived in the existing literature. We shall show that the solutions to the optimization problems in Eq (2) and Eq (3), denoted as $e_{1}^{*}(\cdot),\ldots,e_{T}^{*}(\cdot)$, from a sequence of probabilities that vary across different stages. In comparison, the classical Neyman allocation rule for power maximization:
${e}_{t}^{\mathtt{Neyman}}(a|x)=\frac{\sigma(x,a)}{\sigma(x,1)+\sigma(x,0)},\quad a\in\{0,1\},\quad t=1,\ldots,T,$ (4)
and the optimal failure reduction rule derived in Rosenberger et al. (2001):
${e}_{t}^{\mathtt{FR}}(a|x)=\frac{\sqrt{\mu(x,a)}}{\sqrt{\mu(x,1)}+\sqrt{\mu(x,0)}},\quad a\in\{0,1\},\quad t=1,\ldots,T,$ (5)
are both fixed probabilities that do not differ across stages. Both allocation methods ignore the delay mechanism. Therefore, when response delay plays a significant role, assigning treatments based on existing optimal rules no longer guarantees achieving the desired experimental goals.
Figure 2: Comparison between classical optimal rule and delay-adjusted optimal rule for two design objectives in CARA: (A)-(B) power maximization and (C)-(D) reducing the total number of failures.
We illustrate this point more concretely in a simulation study, where model parameters are calibrated from our synthetic case study and are applied to the oracle optimization problems in Eq (2) and (3). We refer to the detailed data-generating process for this simulation design in Section 5. First, we solve for the oracle treatment allocation with the design objective of maximizing power, as defined in Eq (2). The resulting oracle treatment allocations for each Stage $t$ and the Neyman allocation are presented in Panels (A) and (B) for covariate strata $X=0$ and $X=1$, respectively. The corresponding values of the objective function in Eq (2) under the two different allocations are labeled as $\mathsf{V}_{\text{Delay}}$ and $\mathsf{V}_{\text{Neyman}}$. Second, we solve for the oracle treatment allocation with the design objective of minimizing the total expected number of failures, as in Eq (3). The oracle treatment allocation and the ethical allocation across Stage $t$ are shown in Panels (C) and (D) for the covariate strata $X=0$ and $X=1$, respectively. The percentage of failure under the two allocation strategies is labeled in the figures.
As illustrated in Figure 2(A), the optimal treatment allocation accounting for the response delay (i.e., the solution to Eq (2)) varies across stages and deviates significantly from Neyman allocation. Furthermore, the optimal treatment allocations result in an ATE estimator with reduced variance (e.g., $\mathsf{V}_{\text{Delay}}=1.62$ and $\mathsf{V}_{\text{Neyman}}=2.95$ for the strata with $X=0$). As illustrated in Figure 2(B), a similar observation applies to ethical designs, where the optimal treatment allocation tailored to the ethical design objective (i.e., the solution to Eq (3)) differs from the ethical design rule in Eq (5), and the delay-adjusted optimal allocation achieves a lower failure rate.
Before concluding this section, we note that under appropriate conditions, we will show that the optimal treatment allocation rules $e_{1}^{*}(\cdot),\ldots,e_{T}^{*}(\cdot)$ are unique (Theorem 4). This uniqueness lays the foundation for establishing the statistical guarantees of our proposed design. Moreover, it demonstrates the necessity of accounting for delay mechanisms to achieve the desired experimental design goals.
3 Fully forward-looking CARA with delayed outcomes
The results provided in the previous section assume prior knowledge of the delay mechanism, and therefore, the discussed optimal treatment allocation corresponds to an infeasible oracle setting. In practice, the delay mechanism is rarely known and must be estimated as data are sequentially accumulated to optimize the experimental objective. In this section, we propose two delay-informed practical CARA designs tailored to two previously discussed experimental goals: improving statistical power (Tymofyeyev et al., 2007) and minimizing the expected number of failures while meeting a minimum power requirement for binary responses (Rosenberger et al., 2001). We also provide tools for valid statistical inference after the experiment is finished.
As with any CARA designs, at the first stage of the experiment, we lack information about the participants, so treatments are assigned randomly with a pre-fixed probability. In subsequent stages, our proposed fully forward-looking CARA design incorporates two key components: (1) estimating past delay mechanisms from collected data while anticipating future delays, and (2) using these estimates to formulate a fully forward-looking optimization problem that determines optimal treatment assignments for the next stage. An overview of our proposed fully forward-looking CARA can be found in Figure 3. We introduce these two components in the following paragraphs.
To save notation, we consider a slightly simplified setting in this section that there are only two arms, that is, $a\in\mathcal{A}=\{0,1\}$. We remark that our design strategies can be easily extended to a more general multi-valued treatment setting, with the cost of additional notation. In Section 4, we will establish properties of our proposed design methods and the treatment effect estimators under the general setting, accommodating multiple treatment arms.
Figure 3: Forward looking CARA design illustration.
At the end of each stage $t\geq 2$, we estimate (and update) the observed delay probabilities as
$\hat{\rho}_{t}(d|x,a)=\sum_{\ell=0}^{d}\hat{\mathbb{P}}[D_{i}=\ell|X_{i}=x,A_{i}=a]=\sum_{\ell=0}^{d}\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}=\ell,S_{i}\leq t-\ell)}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,S_{i}\leq t-\ell)}},\quad d\leq t-1,$
for $a=0,1$. Notice that the above estimator is only defined for $d\leq t-1$. To gain some intuition, consider a concrete example for $t=3$. At the end of the third stage, data collected from participants enrolled in the first stage will be informative about the delay mechanism $\rho(0)$, $\rho(1)$, and $\rho(2)$. Among those whose outcome information is still missing at the end of stage 3, it is not possible to tell whether such information will become available until the end of stage 4. Following a similar reasoning, participants enrolled in the second stage can help estimate $\rho(0)$ and $\rho(1)$. The main message is that at any stage $t$, the delay mechanism is only estimable up to $d\leq t-1$, and the remaining “tail features” are not estimable until more data is collected. To anticipate these future delay distributions, we propose the following approaches that reflect the experimenter’s expectations about future delays. For a conservative experimenter, we consider a simple extrapolation approach:
$\hat{\rho}_{t}(d|x,a)=\hat{\rho}_{t}(t-1|x,a),\quad d\geq t.$
This extrapolation essentially assumes that all future delay probabilities will remain at the last estimated value. For an optimistic experimenter, we consider an extrapolation that postulates all missing outcomes will be available after one more experimental stage:
$\hat{\rho}_{t}(d|x,a)=1,\quad d\geq t.$
Lastly, a neutral experimenter might gradually adjust the estimated delay probabilities from the last observed value to immediate response by interpolating between $\hat{\rho}_{t}(t-1|x,a)$ and 1 for $d=t,\dots,T-1$. These estimation strategies enable experimenters to incorporate future delay probabilities into our design, accommodating different perspectives on how delays may evolve. An illustration of different delay mechanism estimation strategies is plotted in Figure 4. Although it is generally impossible to rank the three extrapolation methods, we show through Monte Carlo experiments that they all deliver valid inference for the desired causal effects and more effectively achieves the design objectives compared to classical approaches (such as complete randomization). To avoid unnecessary length discussions, we focus on the conservative approach in our theoretical investigation and show that the optimized treatment allocation is consistent (Theorem 4).
Figure 4: Comparison of different experimenter perspectives on delays and their corresponding estimation on the delay mechanism.
Next, to optimize treatment allocation for stage $t+1$ while accounting for response delays, we introduce a fully forward-looking estimation of the variance term $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ at the current stage $t$. Our proposed estimate incorporates past allocations up to Stage $t$, future allocations from stages $t+1$ to $T$, and both past and future delay mechanisms. It is designed to anticipate the impact of future allocations and outcome delays on the variance of the final treatment effect estimate after the experiment is finished. Our estimate is of the following form:
$\hat{\mathsf{V}}^{\text{forward}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Bigg{[}\frac{\hat{\sigma}_{t}^{2}(x,1)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)\hat{e}^{*}_{\ell}(1|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)e_{\ell}(x)}$
$\qquad\quad+\frac{\hat{\sigma}_{t}^{2}(x,0)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)\hat{e}^{*}_{\ell}(0|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)(1-e_{\ell}(x))}\Bigg{]}.$
In this expression, $\hat{p}_{t}(x)$ represents the estimated proportion of participants with covariate $x$ up to stage $t$, calculated as
$\hat{p}_{t}(x)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,S_{i}\leq t)}}{N_{t}},$
where recall that $n_{\ell}$ is the number of participants enrolled at stage $\ell$, $N_{t}$ is the total number of participants up to stage $t$, and $N=N_{T}$. The term $\hat{r}_{\ell}=n_{\ell}/N$ denotes the proportion of participants enrolled at stage $\ell$. The estimated conditional variance of the outcome $Y(a)$ given covariate $x$ and treatment $a$, up to stage $t$, is denoted by $\hat{\sigma}_{t}^{2}(x,a)$ and is calculated as
$\hat{\sigma}_{t}^{2}(x,a)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}(Y_{i}-\hat{\mu}_{t}(x,a))^{2}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}},\quad\hat{\mu}_{t}(x,a)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}Y_{i}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}}.$
While the expressions may seem complicated, they are simply the sample mean and sample variance of the outcome variable, restricted to the subsample for group $x$ receiving treatment $a$.
Using this fully forward-looking variance estimate, we can then determine an effective treatment assignment for stage $t+1$ that accounts for how future treatment allocations and response delays from stages $t+1$ onward affect the design objectives at the end of the experiment. For the first design objective of maximizing power, we solve the following optimization problem to find the treatment allocation probability for stage $t+1$:
$(\tilde{e}_{t+1}(\cdot),\dots,\tilde{e}_{T}(\cdot))=\operatorname*{argmin}_{e_{t+1},\ldots,e_{T}}\left\{\hat{\mathsf{V}}^{\text{forward}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot)):\ \delta\leq e_{\ell}(\cdot)\leq 1-\delta,\ \ell=t+1,\ldots,T\right\}.$ (6)
We then set $\hat{e}^{*}_{t+1}(\cdot)=\tilde{e}_{t+1}(\cdot)$.
For the second design objective of reducing failures under a power constraint, we define a similar fully forward-looking estimate of the expected proportion of total failures as
$\hat{\mathsf{P}}^{\text{forward}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)=$ $\sum_{\ell=1}^{t}\hat{r}_{\ell}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\left[\hat{e}^{*}_{\ell}(1|x)\left(1-\hat{\mu}_{t}(x,1)\right)+\hat{e}^{*}_{\ell}(0|x)\left(1-\hat{\mu}_{t}(x,0)\right)\right]$
$+\sum_{\ell=t+1}^{T}r_{l}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\left[e_{t}(1|x)\left(1-\hat{\mu}_{t}(x,1)\right)+e_{t}(0|x)\left(1-\hat{\mu}_{t}(x,0)\right)\right].$
We denote the solution to the following optimization problem as $(\tilde{e}_{t+1}(\cdot),\dots,\tilde{e}_{T}(\cdot))$:
$\min_{e_{t+1},\ldots,e_{T}}$ $\ \hat{\mathsf{P}}^{\text{forward}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)$ (7)
subject to $\ \delta\leq e_{\ell}(\cdot)\leq 1-\delta,\quad\text{for }\ell=t+1,\ldots,T$
$\ \hat{\mathsf{V}}^{\text{forward}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)+\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\big{(}\hat{\tau}_{t}(x)-\hat{\tau}_{t}\big{)}^{2}\leq C,$
where $\hat{\tau}_{t}(x)=\hat{\mu}_{t}(x,1)-\hat{\mu}_{t}(x,0)$ and $\hat{\tau}_{t}=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\hat{\tau}_{t}(x)$. We then again set $\hat{e}^{*}_{t+1}(\cdot)=\tilde{e}_{t+1}(\cdot)$.
To conclude our design strategy, as both optimization problems in Eqs  (6) and (7) are fully forward-looking, our proposed design strategies aim to proactively adjust the allocation strategies in future stages, ensuring that the data collected accounts for the future impacts due to outcome delays. This approach contrasts with classical CARA designs that optimize trial objectives solely using historical information. At the end of the last stage $T$, the average treatment effect estimator is constructed with
$\hat{\tau}_{T}=\sum_{x\in\mathcal{X}}\hat{p}_{T}(x)\hat{\tau}_{T}(x)=\sum_{x\in\mathcal{X}}\hat{p}_{T}(x)\Big{(}\hat{\mu}_{T}(x,1)-\hat{\mu}_{T}(x,0)\Big{)}.$
We then provide a variance estimator, thus enabling valid statistical inference:
$\hat{\mathsf{V}}_{T}$ $=\sum_{x\in\mathcal{X}}\hat{p}_{T}(x)\Bigg{[}\frac{\hat{\sigma}_{T}^{2}(x,1)}{\frac{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=1,D_{i}+S_{i}\leq T)}}{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x)}}}+\frac{\hat{\sigma}_{T}^{2}(x,0)}{\frac{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=0,D_{i}+S_{i}\leq T)}}{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x)}}}+\big{(}\hat{\tau}_{T}(x)-\hat{\tau}\big{)}^{2}\Bigg{]}.$
In the following sections, we will provide theoretical guarantees and simulation results to demonstrate the advantages of our proposed fully forward-looking CARA designs in the presence of delayed outcomes.
4 Theoretical investigation
This section provides theoretical justifications for the proposed forward-looking CARA design. To begin, we present the semiparametric efficiency bound for estimating average treatment effects in the presence of response delay in multi-stage adaptive experiments. We then develop a general result on the consistency and asymptotic normality of the estimated average treatment effect.
Our theoretical investigation then proceeds by examining the specific setting considered in Section 3, where we extrapolate the unknown delay mechanism in each stage using a conservative rule. Encouragingly, we are able to justify Condition 1 in this setting under mild primitive assumptions. Finally, we observe that in an experiment with only finitely many stages, it is generally not possible to achieve the oracle-optimal allocation, as this would require perfect knowledge of the delay mechanism. Therefore, the last part of our theoretical investigation addresses the efficiency loss resulting from extrapolating the delay mechanism in our design method. Under conservative extrapolation, we show that the optimized asymptotic variance is guaranteed to weakly decrease as the experiment progresses.
Our main results are developed under the general setting of multi-valued treatments, that is, there can be more than two treatment arms. Therefore, we will adopt a more general notation for the estimated treatment effects. As before, let $\hat{\mu}_{T}(x,a)$ denote the estimated mean potential outcome for treatment arm $a$ and subgroup $x$. Then the estimated subgroup treatment effect is $\hat{\tau}_{T}(x,a^{\prime},a)=\hat{\mu}_{T}(x,a^{\prime})-\hat{\mu}_{T}(x,a)$, and the estimated average treatment effect is $\hat{\tau}_{T}(a^{\prime},a)=\sum_{x\in\mathcal{X}}\hat{p}_{T}(x)\hat{\tau}_{T}(x,a^{\prime},a)$.
4.1 Notation, assumptions and semiparametric efficiency bound
We begin by recalling our adaptive experiment setting: the experiment consists of $T$ stages, and in each stage $t=1,2,\dots,T$, $n_{t}$ participants are enrolled. Let $N_{t}=\sum_{s=1}^{t}n_{s}$ represent the cumulative sample size at the end of stage $t$, with $N=N_{T}$ as the total sample size. We use the notation $S_{i}=t$ to indicate that individual $i$ is enrolled in stage $t$.
For each individual, we observe some baseline covariates, $X_{i}\in\mathcal{X}$. The observed outcome variable is denoted by $Y_{i}$, and it is related to the potential outcomes through the relation $Y_{i}=\sum_{a\in\mathcal{A}}\mathds{1}_{(A_{i}=a)}Y_{i}(a)$, where $A_{i}$ denotes the actual treatment assigned. Our first assumption imposes standard regularity conditions on the covariates and the potential outcomes.
Assumption 1 (Covariates and potential outcomes).
(i) The covariates and potential outcomes, $\{X_{i},\{Y_{i}(a):a\in\mathcal{A}\}\}$, are independently and identically distributed across $i=1,2,\dots,N$. (ii) The covariates have finite support, that is, $|\mathcal{X}|<\infty$. Let $p(x)=\mathbb{P}[X_{i}=x]$, then $\min_{x\in\mathcal{X}}p(x)>0$. (iii) The potential outcomes have finite fourth moments: $\max_{x\in\mathcal{X}}\max_{a\in\mathcal{A}}\mu_{4}(x,a)<\infty$.
The next assumption addresses the delay mechanism. Recall that the delay is represented by the random variable $D_{i}$, which indicates the number of stages after which the experimenter can observe $Y_{i}$. By our definition, $Y_{i}$ is observed at the end of stage $t$ if and only if $D_{i}+S_{i}\leq t$.
Assumption 2 (Delay mechanism).
[…]
Finally, as we consider multi-stage experiment settings, we are in an asymptotic regime in which $T$ is fixed, and $n_{t}\to\infty$.
Assumption 3 (Asymptotic regime).
The total number of stages, $T$, is fixed, and for all $t$, $n_{t}/N\to r_{t}>0$.
We are now ready to present the semiparametric efficiency bound for average treatment effect estimation in the oracle setting, where there is perfect knowledge of the delay mechanism, and treatments are randomly assigned conditional on the covariates. This efficiency bound serves as the foundation and starting point for our proposed adaptive experimental design, as our algorithm minimizes a feasible version of the efficiency bound, constructed using accrued information.
Theorem 1 (Semiparametric efficiency bound).
Let Assumptions 13 hold. In addition, assume (i) $S_{i}$ are independently and identically distributed with $\mathbb{P}[S_{i}=t]=r_{t}$; (ii) the treatment assignment probabilities, $e_{t}(a|x)=\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t]$, are bounded away from 0 and 1. Then the efficient influence function for estimating $\mu_{1}(a)$ is
$\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a)$ $=\frac{\mathds{1}_{(A_{i}=a,D_{i}+S_{i}\leq T)}}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|X_{i})\rho(T-t|X_{i},a)}\Big{(}Y_{i}(a)-\mu_{1}(X_{i},a)\Big{)}+\mu_{1}(X_{i},a)-\mu_{1}(a).$
In addition, the efficiency influence function for estimating $\tau(a^{\prime},a)$ is
$\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a^{\prime},a)\ =\ \psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a^{\prime})-\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a)$
$=$ $\frac{\mathds{1}_{(A_{i}=a^{\prime},D_{i}+S_{i}\leq T)}}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a^{\prime}|X_{i})\rho(T-t|X_{i},a^{\prime})}\Big{(}Y_{i}(a^{\prime})-\mu_{1}(X_{i},a^{\prime})\Big{)}$
$\qquad\qquad-\frac{\mathds{1}_{(A_{i}=a,D_{i}+S_{i}\leq T)}}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|X_{i})\rho(T-t|X_{i},a)}\Big{(}Y_{i}(a)-\mu_{1}(X_{i},a)\Big{)}+\tau(X_{i},a^{\prime},a)-\tau(a^{\prime},a),$
leading to the semiparametric efficiency bound:
$\mathsf{V}=\sum_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,a^{\prime})}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a^{\prime}|x)\rho(T-t|x,a^{\prime})}$ $+\frac{\sigma^{2}(x,a)}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}+\big{(}\tau(x,a^{\prime},a)-\tau(a^{\prime},a)\big{)}^{2}\Bigg{]}.$
We note that while the previous theorem is developed in the context of discrete covariates (with the efficiency bound expressed as a summation over the probability mass function), it can be generalized to accommodate, for example, continuously distributed covariates (c.f., equation 1).
4.2 Asymptotic properties of treatment effects estimates
In this subsection, we demonstrate that the treatment effect estimator is consistent and admits an asymptotically normal distribution. The main conclusions of this section rely on a high-level condition regarding the consistency of the optimized treatment allocation rule, making these conclusions applicable to a wide range of CARA and a broad class of frequentist adaptive experimental design settings (Hu and Zhang, 2004; Baldi Antognini and Zagoraiou, 2015; Hu et al., 2015). In line with existing literature, the consistency of optimized treatment allocation rules is often needed to establish statistical guarantees of various adaptive designs. We recall that the treatment assignment probabilities are denoted by $\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t,\mathcal{H}_{t-1}]=\hat{e}^{*}_{t}(a|x)\in\mathcal{H}_{t-1}$, which are allowed to depend on historical data.
We begin with a high-level condition, which will be verified in the next subsection for our proposed forward-looking CARA design.
Condition 1 (Convergence of the optimized treatment allocation).
There exists nonrandom $e_{t}(\cdot|\cdot)$, such that $\hat{e}^{*}_{t}(a|x)=e_{t}(a|x)+o_{\mathrm{p}}(1)$ for all $a\in\mathcal{A}$ and all $x\in\mathcal{X}$.
We then present a useful lemma, which characterizes an asymptotically linear representation of the estimated mean potential outcomes.
Lemma 1 (Asymptotically linear representation for estimated mean potential outcomes).
Let Assumptions 13 hold. Then under Condition 1,
$\sqrt{N}\big{(}\hat{\mu}_{1,T}(x,a)-\mu_{1}(x,a)\big{)}$
$\quad=\frac{1}{p(x)\sum_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq T)}\big{(}Y_{i}-\mu_{1}(x,a)\big{)}+o_{\mathrm{p}}(1).$
The next theorem characterizes the asymptotic normality of both the estimated subgroup treatment effects and the estimated average treatment effect.
Theorem 2 (Asymptotic normality of estimated subgroup treatment effects).
Let Assumptions 13 hold. Then under Condition 1,
$\sqrt{N}\big{(}\hat{\tau}_{T}(x,a^{\prime},a)-\tau(x,a^{\prime},a)\big{)}\overset{\mathrm{d}}{\to}\mathcal{N}\Big{(}0,\frac{1}{p(x)}\Big{(}\frac{\sigma^{2}(x,a^{\prime})}{\sum_{t=1}^{T}r_{t}e_{t}(a^{\prime}|x)\rho(T-t|x,a^{\prime})}+\frac{\sigma^{2}(x,a)}{\sum_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}\Big{)}\Big{)},$
and
$\sqrt{N}\big{(}\hat{\tau}_{T}(a^{\prime},a)-\tau(a^{\prime},a)\big{)}\overset{\mathrm{d}}{\to}\mathcal{N}(0,\mathsf{V}),$
where $\mathsf{V}$ is defined in Theorem 1.
Before discussing statistical inference, we compare this theorem to the semiparametric efficiency bound established in the previous subsection. Specifically, while the two results, Theorems 1 and 2, yield the same variance formula, they are conceptually very different. The semiparametric efficiency bound is derived under a fixed treatment assignment regime, which we then use to guide our experimental design algorithm. Specially, our algorithm aims to minimize this efficiency bound. However, from an ex ante perspective, it may be unclear why this is a useful exercise, as it is not immediately apparent how to achieve this efficiency bound. Encouragingly, Theorem 2 helps “close the loop” by showing that, under mild regularity conditions, the average treatment effect estimator achieves the efficiency bound, thereby justifying our design objective. To conclude this subsection, the result below establishes valid statistical inference, where the variance estimator, $\hat{\mathsf{V}}_{T}$, is defined at the end of Section 3.
Theorem 3 (Statistical inference).
Let Assumptions 13 hold. Then under Condition 1, $\hat{\mathsf{V}}_{T}={\mathsf{V}}_{T}+o_{\mathrm{p}}(1)$.
4.3 Convergence of the optimized treatment allocation
Previously we have shown that the treatment effect estimators are consistent and asymptotically normally distributed. These results build on the high-level condition that the empirical treatment allocation, $\hat{e}_{t}^{*}(\cdot)$, will converge in large samples (Condition 1). In this subsection, we further investigate the theoretical properties of the proposed experimental design—specifically, solved from (6)—and show that it satisfies Condition 1, thereby tying all loose ends. To give a road map, we will first show that the asymptotic analogs of our design algorithm have unique solutions. In this process, we will carefully distinguish the oracle problem of minimizing the semiparametric efficiency bound (which requires perfect knowledge of the delay mechanism) from the problem of sequential optimization using only accrued knowledge (which requires extrapolating the delay mechanism). Next, we show that the empirically optimized treatment allocation converges to some large-sample limit. To conclude this subsection, we will present a bound on efficiency loss due to extrapolating the delay mechanism.
In our theoretical investigation, we will focus on the conservative method for estimating the delay mechanism. This allows us to provide simple and easy-to-interpret conditions, hence avoiding lengthy discussions. Despite our focus on one particular approach in this subsection, we show in our simulations that all three methods (conservative, optimistic, and neutral) deliver valid inference for the desired causal effects and more effectively achieves the design objectives compared to classical approaches.
To start, we recall that $\mathsf{V}=\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ denotes the efficiency bound. Its minimizers will be denoted by
$\{e_{t}^{*}(\cdot),1\leq t\leq T\}=\operatorname*{argmin}_{e_{1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot)),\ \text{subject to }e_{t}(a|\cdot),e_{t}(a^{\prime}|\cdot)\geq\delta.$
As the semiparametric efficiency bound is specific to two treatment values, $a$ and $a^{\prime}$, we impose the constraints that the treatment probabilities in these two arms are bounded from below. However, it is worth mentioning that one can easily incorporate additional constraints.
As we have discussed, the oracle optimal solution $\{e_{t}^{*}(\cdot),1\leq t\leq T\}$ is not achievable in general, as it builds on perfect knowledge of the delay mechanism. In any realistic experimental setting, however, this information can only be learned sequentially. In fact, even we set $n_{t}=\infty$, part of the delay mechanism can still remain unknown for any $t$: to be even more precise, one typically only learn $\rho(d|\cdot)$ for $d<t$ using information at the end of stage $t$. For this reason, we have introduced feasible and concrete ways to extrapolate the delay mechanism. Following our discussion in Section 2, we define
$\rho^{\dagger}_{t}(d|\cdot)$ $=\min\big{\{}\rho(d|\cdot),\rho(t-1|\cdot)\big{\}}\ \text{for $d=1,2,\dots$}.$
Alternatively, we write $\rho^{\dagger}_{t}(d|\cdot)=\rho(d|\cdot)$ for $d<t$, and $\rho^{\dagger}_{t}(d|\cdot)=\rho(t-1|\cdot)$ for $d=t,t+1,\dots$. It is helpful to compare this definition with Assumption 2. There, we assumed that the delay mechanism is time-homogeneous, so that $\rho(\cdot)$ is not indexed by $t$. On the other hand, $\rho^{\dagger}_{t}(\cdot)$ arises due to our imperfect knowledge about the delay mechanism and the conservative extrapolation employed, and as a result it will depend on a specific stage. It is not the true delay mechanism, but it is the asymptotic analogue of $\hat{\rho}_{t}(\cdot)$, which we used in our optimization algorithm.
To study convergence property of $\hat{e}^{*}_{t}(\cdot)$, we will make a definition first. Let $e^{\dagger}_{1}(a|\cdot)=e^{\dagger}_{1}(a^{\prime}|\cdot)=1/2$. Then for $t=2,3,\dots,T$, define recursively that
$e^{\dagger}_{t}(\cdot)=\operatorname*{argmin}_{e_{t}(\cdot)}\min_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t}^{\dagger}(e_{t}(\cdot),e_{t+1}(\cdot),\dots,e_{T}(\cdot)),\ \text{subject to }e_{t}(a|\cdot),e_{t}(a^{\prime}|\cdot)\geq\delta,$
where
$\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,a^{\prime})}{\sum\limits_{\ell=1}^{t}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a^{\prime})e_{\ell}^{\dagger}(a^{\prime}|x)+\sum\limits_{\ell=t+1}^{T}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a^{\prime})e_{\ell}(a^{\prime}|x)}$
$+\frac{\sigma^{2}(x,a)}{\sum\limits_{\ell=1}^{t}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a)e_{\ell}^{\dagger}(a|x)+\sum\limits_{\ell=t+1}^{T}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a)e_{\ell}(a|x)}\Bigg{]}.$
Our first main result is to show that both $e^{*}_{t}(\cdot)$ and $e^{\dagger}_{t}(\cdot)$ are unique. To this end, the following assumption.
Assumption 4 (Variation in delay mechanism).
For any $d\neq d^{\prime}$, ${\rho(d|x,a^{\prime})}/{\rho(d|x,a)}\neq{\rho(d^{\prime}|x,a^{\prime})}/{\rho(d^{\prime}|x,a)}$.
We remark that this assumptions helps establish the uniqueness of $e^{*}_{t}(\cdot)$. However, it is still possible to have ${\rho_{t}^{\dagger}(d|x,a^{\prime})}/{\rho_{t}^{\dagger}(d|x,a)}={\rho_{t}^{\dagger}(d^{\prime}|x,a^{\prime})}/{\rho_{t}^{\dagger}(d^{\prime}|x,a)}$. Fortunately, this is easily resolved by a tie-breaking rule. See the Supplementary Materials for a specific recommendation and additional discussions.
Theorem 4 (Consistent treatment allocation).
Let Assumptions 14 hold. Then both $e_{t}^{*}(\cdot)$ and $e_{t}^{\dagger}(\cdot)$ are unique. In addition, Condition 1 holds with $\hat{e}_{t}^{*}(\cdot)\overset{\mathrm{p}}{\to}e_{t}^{\dagger}(\cdot)$.
Before closing this section, we provide two insights regarding the variance minimization problem. The first result suggests that as more information about the delay mechanism becomes available in the adaptive experiment, the optimized asymptotic variance never increases. While this result may seem quite natural, we remark that it does rely on the specific extrapolation we employed for imputing the delay mechanism, and to be specific, the conservative method we discussed earlier.
The second result provides a bound on efficiency loss due to imperfect knowledge about the delay mechanism. Collectively, these two results provide theoretical guarantee to our proposed adaptive experimental design.
Theorem 5 (Optimized variance and bound on efficiency loss).
Let Assumptions 14 hold. Define $\mathsf{V}^{*}$ as the minimized oracle asymptotic variance. Then for all $t=1,2,\dots,T-1$,
$\min_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))\ \geq$ $\min_{e_{t+2}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t+1}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))\ \geq\ \mathsf{V}^{*}.$
Denote by $\mathsf{V}$ the asymptotic variance of the estimated average treatment effect. Then $\mathsf{V}\leq\mathsf{V}^{*}+O(1/T)$.
5 Synthetic case study
To evaluate the performance of the proposed forward-looking CARA design, we conduct a synthetic case study based on the randomized experiment reported in Fahey et al. (2020). Through this synthetic case study, we first demonstrate that both of the two forward-looking CARA designs (Section 3) deliver valid statistical inference (Table 1), supporting the theoretical result established in Theorem 2. Second, numerical evidence suggests that our fully forward-looking CARA design with the power maximization objective achieves a higher estimation efficiency compared to both the Neyman allocation and the complete randomization design in the presence of delayed responses (Figures 5 and 6), further validating the theoretical results in Theorem 5. Finally, compared to both the ethical design (Rosenberger et al., 2001) and complete randomization, our fully forward-looking CARA design can provide greater failure reduction, demonstrating its effectiveness in improving participant welfare while maintaining statistical power (Figure 5).
The original experiment conducted by Fahey et al. (2020) aims to evaluate the effectiveness of cash transfers on retention in care among people living with HIV (PLHIV), using viral load as the primary response variable. The experiment participants were adults aged 18 years or older diagnosed with HIV, who were randomized to receive cash transfers (treatment arm) or no cash transfers (control arm). The response variable, viral load, was assessed six months after enrollment. Although the original study was not conducted in an adaptive fashion, adaptation can be applied since the participants were enrolled sequentially. However, a challenge is that the primary responses of the participants were collected with delays, with some responses collected longer than six months after the treatment assignment. As illustrated in Example 1, the empirically observed delay mechanism exhibits significant heterogeneity across arms and covariate strata. Our fully forward-looking CARA design can be utilized in such scenarios to help the experiment more effectively achieve its design goals. In this synthetic case study, we conduct a simulation experiment in which the underlying data-generating process — including the response delay mechanism, covariate distribution, and the (conditional) potential outcomes distributions — is calibrated with estimates from the (Fahey et al., 2020) study. We set $n_{t}=100$ for $t=1,\ldots,T$, and $T$ ranges in $\{4,6,8\}$.
The primary response variable is the viral load measured at the end of each stage. We consider two different transformations of the original viral load measurements, which shall be illustrated in the two simulation setups below. Delays in collecting the response variable are observed, and as discussed in Example 1, the delay mechanism depends on the arm and the biological sex of the participant. We consider two treatment arms: the control arm $A_{i}=0$, defined as “received no cash transfer”, and the treatment arm, $A_{i}=1$, defined as “received cash transfer.” Biological sex is included as the covariate $X_{i}$, where $X_{i}=1$ and 00 represent “male” and “female,” respectively. From the original experimental data, the distribution of biological sex consists of 36% male and 64% female, which we adopt as the data-generating process of $X_{i}$ in our Monte Carlo exercise. Specifically, we generate the covariate $X_{i}$ and treatment $A_{i}$ as follows:
$X_{i}$ $\sim\text{Bernoulli}(0.36),\ i=1,\ldots,N,$
$A_{i}|X_{i}=x,S_{i}=t$ $\sim\text{Bernoulli}(e_{t}(1|x)),\ t=1,\ldots,T,$
where $e_{t}(1|x)$ entails different designs we aim to compare. In what follows, we shall illustrate the design strategies in comparison under two sets of simulation setups tailored to each design objective:
Setup 1: Power maximization. For the power maximization design objective in Eq (2), we compare four designs: (1) the fully forward-looking CARA design for power maximization (Section 3); (2) the fully forward-looking CARA design further enhanced by the doubly adaptive biased coin (DBCD) method (Hu and Zhang, 2004; Tymofyeyev et al., 2007); (3) complete randomization; (4) treatment allocation based on Neyman allocation. Furthermore, in HIV research, it is standard practice to apply a log transformation to raw viral load measurements, given their wide range. Consequently, the response variable in this simulation setup is defined as: $\log(\texttt{viral load})$. We generate the response variable as
$Y_{i}|A_{i}=a,X_{i}=x\sim\mathcal{N}\left(\mu(x,a),\ \sigma^{2}(x,a)\right)$
to mimic the distribution of $\log(\texttt{viral load})$ in the original data, where for women $\mu(0,1)=2.50$, $\mu(0,0)=2.98$, $\sigma(0,1)=0.36$, $\sigma(0,0)=2.06$, and for men $\mu(1,1)=2.47$, $\mu(1,0)=2.72$, $\sigma(1,1)=0.82$, and $\sigma(1,0)=0.31$. Lastly, to evaluate the power of different designs under the power maximization objective, we vary the magnitude of ATE in our data-generating process by perturbing $\mu(1,0)$, such that the ATE $\tau(1,0)=0.33+\delta$, where $\delta\in\{-0,33,\ldots,0,\ldots,0.33\}$.
Setup 2: Failure reduction. For the failure reduction design objective in Eq (3), we compare four designs: (1) the proposed fully forward-looking CARA design for failure reduction; (2) the proposed design enhanced by the DBCD method (Hu and Zhang, 2004); (3) the ethical design introduced by Rosenberger et al. (2001); (4) complete randomization. In this simulation setup, since the goal is to reduce the percentage of failures, we transform the original viral load measurements into a binary response variable, indicating whether the participant has achieved viral suppression. Specifically, if a participant’s viral load is less than $1,000$ copies per mL, we set $Y_{i}=1$, representing viral suppression. Otherwise, $Y_{i}$ is set to 00 (no viral suppression). We generate the binary response variable as
$Y_{i}|A_{i}=a,X_{i}=x\sim\text{Bernoulli}(\mu(x,a)),\ i=1,\ldots,N,$
where $\mu(0,1)=0.78$ and $\mu(0,0)=0.57$ represent the expected probabilities of viral suppression for females in the treatment and control arms, respectively. Similarly, $\mu(1,1)=0.84$ and $\mu(1,0)=0.63$ represent the expected probabilities of viral suppression probabilities for males in the treatment and control arms, respectively. Lastly, to evaluate the power of different designs under the failure reduction objective, we vary the magnitude of ATE in our data-generating process. For Setup 2, we vary the ATE value by perturbing $\mu(1,0)$, such that the ATE $\tau(1,0)=0.21+\delta$, where $\delta\in\{-0,21,\ldots,0,\ldots,0.21\}$.
In our synthetic case study, we incorporate both arm- and covariate-specific outcome delays following empirical estimates obtained in the original experimental data. We generate the delay variable $D_{i}$ from multinomial distribution:
$D_{i}|A_{i}=1,X_{i}=1$ $\sim\text{Multinomial}\left(1,\bm{\rho}(\cdot|1,1)\right),\ D_{i}|A_{i}=1,X_{i}=0\sim\text{Multinomial}\left(1,\bm{\rho}(\cdot|1,0)\right),$
$D_{i}|A_{i}=0,X_{i}=1$ $\sim\text{Multinomial}\left(1,\bm{\rho}(\cdot|0,1)\right),\ D_{i}|A_{i}=0,X_{i}=0\sim\text{Multinomial}(1,\bm{\rho}(\cdot|0,0)),$
where the first row represents the response delay mechanism for males and females under the treatment arm, while the second row represents the delay mechanism for males and females under the control arm. We slightly abuse the notation to define $\bm{\rho}(\cdot|a,x):=(\mathbb{P}(D_{i}=0|a,x),\ldots,\mathbb{P}(D_{i}=T-1|a,x))^{\intercal}$, where each element represents the probability that the response of subject $i$ is delayed by exactly $d$ stages, for $d=0,\ldots,T-1$. When $T=4$, the parameter values for the delay mechanisms are as follows: $\bm{\rho}(\cdot|1,1)=(0.64,0.18,0.07,0.03)^{\intercal}$, $\bm{\rho}(\cdot|1,0)=(0.63,0.18,0.05,0.02)^{\intercal}$, $\bm{\rho}(\cdot|0,1)=(0.55,0.23,0.10,0.02)^{\intercal}$, and $\bm{\rho}(\cdot|0,0)=(0.54,0.11,0.21,0.01)^{\intercal}$. We defer the details of the delay mechanism generation for $T=6$ and $T=8$ to the Supplementary Materials. To generate delayed responses, at each Stage $t$, we first generate responses for all $n_{t}$ subjects. Based on the delay mechanism, we then randomly select a subset of subjects with $(A_{i}=a,X_{i}=x)$ and label their responses as NA, indicating that their responses will be delayed to be observed.
For both simulation setups, we generate 500 Monte Carlo samples. To initialize the experiment, we set the treatment allocation to be $1/2$ in each covariate stratum in Stage 1 for all the designs. For Setup 1, we shall compare the variance of the ATE estimator under different design strategies in Figure 5 (A) – (C). For Setup 2, we shall compare the failure reduction rate under different design strategies in Figure 5 (D) – (E). For Setup 1 and Setup 2, we consider the conservative experimenter’s expectations about future delays. We defer the simulation results for neutral and optimistic experimenter’s expectations to Supplementary Materials.
For both setups, we also present the coverage probabilities (Table 1) and the power curve (Figure 6) to evaluate the statistical validity.
Figure 5: The variance of the ATE estimator under the power maximization design objective (A) – (C) and the failure reduction rate under the failure reduction design objective (D) – (F) across different design strategies.
Figure 6: The power curves under the (A) power maximization design objective and the (B) failure reduction design objective across various design strategies.
Design 1: Power Maximization
Delay perspective Proposed CARA design DBCD + Proposed CARA design
Conservative 0.95 (0.01) 0.94 (0.01)
Optimistic 0.94 (0.01) 0.96 (0.01)
Neutral 0.95 (0.01) 0.96 (0.01)
Design 2: Failure Reduction
Delay perspective Proposed CARA design DBCD + Proposed CARA design
Conservative 0.96 (0.01) 0.95 (0.01)
Optimistic 0.95 (0.01) 0.94 (0.01)
Neutral 0.94 (0.01) 0.95 (0.01)
Table 1: Coverage probability comparison across different delay perspectives under two different design objectives.
Figure 5 (A) – (C) shows the variance of the ATE estimator under the power maximization objective across experimental stages for different designs. The results show that the proposed fully forward-looking CARA design and its DBCD-enhanced counterpart exhibit a smaller deviation from the oracle variance (indicated by the dashed line) than the complete randomization design and the Neyman allocation. This suggests that the proposed forward-looking design achieves higher estimation efficiency. Furthermore, Figure 5 (A) – (C) demonstrates that the efficiency of the fully forward-looking CARA design progressively improves as the experimental stages advance. We conjecture that this is because, as we obtain more information about the delay mechanism, the treatment allocation approaches the optimal allocation, leading to more significant changes over time. In contrast, while the Neyman allocation design has a notable improvement in efficiency after Stage 1, its variance stabilizes in subsequent stages. This is because, after the first stage, the Neyman allocation design adjusts for the $1/2$ treatment allocation initialized in Stage 1. Since Neyman allocation does not account for the delay mechanism, its variance quickly stabilizes over time.
Figure 5 (D) – (E) shows a similar trend under the failure reduction design objective. The fully forward-looking CARA design and its DBCD-enhanced counterpart achieve high failure reduction rates. Moreover, while the fully forward-looking CARA design shows a gradual decline in failure reduction as experimental stages increase, the failure reduction rates for the ethical design and the complete randomization design remain stable over time. This suggests that the adaptive nature of the fully forward-looking CARA design facilitates dynamically improved performance across stages. Lastly, our proposed fully forward-looking CARA design enhanced by DBCD overall exhibits a more robust performance. We conjecture that this is because the DBCD design targets the current stage treatment allocation toward the optimal treatment allocation, thereby further improving the empirical performance (i.e., estimation efficiency and failure rate reduction).
As our design is rigorously guided by the semiparametric efficiency theory to improve the estimation efficiency of the average causal effect, we present power curves for correctly detecting the true ATEs in Figure 6. Figure 6 suggests that when the ATE is 0, all considered designs control the type I error rate at the nominal level. Across two design objectives, the proposed forward-looking design overall demonstrates higher power in detecting the treatment effect compared to other designs, further validating the estimation efficiency of our proposed design.
To demonstrate the statistical validity of our proposed designs under different experimenters’ anticipation of delay mechanism, we present the coverage probabilities of the constructed confidence interval correctly covering the true ATE under both design objectives in Table 1. Table 1 shows that both the proposed fully forward-looking CARA design and its DBCD-enhanced design attain the nominal-level coverage probabilities across various delay perspectives. This suggests that, regardless of the perspective on delay when carrying out the experiment, the resulting ATE estimators under the proposed fully forward-looking design provide valid statistical inferences.
In summary, this synthetic case study demonstrates that our proposed fully forward-looking designs provide valid statistical inference in the presence of delayed responses. Additionally, our proposed designs exhibit higher estimation efficiency under the power maximization objective and improved failure reduction rate under the failure reduction objective compared to the competing designs, further validating the theoretical results presented in Section 4.
6 Discussion and concluding remarks
In this manuscript, we have proposed a novel fully forward-looking CARA design to address the challenges posed by delayed outcomes. By accounting for response delay mechanisms to identify optimal treatment allocations, our design achieves substantial improvements in experimental efficiency and participant welfare compared to traditional CARA designs. Through rigorous theoretical development, we derived semiparametric efficiency bounds that account for covariate-dependent delays and demonstrated that the proposed design achieves consistent and asymptotically normal treatment effect estimators. Our simulation studies further validated the practical advantages of this CARA design, showcasing its ability to outperform classical methods like Neyman allocation and complete randomization in terms of statistical efficiency and ethical outcomes. Notably, the forward-looking design dynamically adapts to evolving data, ensuring robust performance across varying experimental stages. Moreover, our work provides practical tools for implementing delay-informed CARA designs, offering pathways for future studies in clinical trials and other adaptive experimental settings.

### v2

@@
-Covariate Adjusted Response Adaptive Design with Delayed Outcomes
+Xinwei Ma Jingshen Wang Waverly Wei
 Abstract
-Covariate-adjusted response adaptive (CARA) designs have gained widespread adoption for their clear benefits in enhancing experimental efficiency and participant welfare.
+Covariate-adjusted response-adaptive (CARA) designs have gained widespread adoption for their clear benefits in enhancing experimental efficiency and participant welfare.
 These designs dynamically adjust treatment allocations during interim analyses based on participant responses and covariates collected during the experiment.
 However, delayed responses can significantly compromise the effectiveness of CARA designs, as they hinder timely adjustments to treatment assignments when certain participant outcomes are not immediately observed.
-In this manuscript, we propose a fully forward-looking CARA design that dynamically updates treatment assignments throughout the experiment as response delay mechanisms are progressively estimated.
-Our design strategy is informed by novel semiparametric efficiency calculations that explicitly account for outcome delays in a multi-stage adaptive experiment.
+In this paper, we propose a fully forward-looking CARA design that dynamically updates treatment assignments throughout the experiment as response delay mechanisms are progressively estimated.
+Our design strategy is informed by novel semiparametric efficiency calculations that explicitly account for outcome delays in a multi-stage setting.
 Through both theoretical investigations and simulation studies, we demonstrate that our proposed design offers a robust solution for handling delayed outcomes in CARA designs, yielding significant improvements in both statistical power and participant welfare.
-Keywords: Causal inference; Delayed outcomes; Frequentist adaptive experimental design; Response adaptive designs
+Keywords: Delayed outcomes; Frequentist adaptive experimental design; Response adaptive designs.
 1 Introduction
-1.1 Motivation and our contribution
-In clinical trials, field experiments in social sciences, and A/B tests, participants often enroll sequentially, and responses to treatments can vary due to individual-specific characteristics.
-In these contexts, covariate-adjusted response-adaptive (CARA) designs have gained increasing attention for their ability to enhance both the efficiency and the ethical or welfare considerations of experimental outcomes (Zhang et al., 2007; Hu et al., 2015; Robertson et al., 2023; Wei et al., 2023, 2024; Zhao, 2023).
-By dynamically adjusting treatment assignments based on accumulated responses and covariates during the experiment, CARA designs can be adopted to optimize treatment allocation to improve statistical power, reduce participant exposure to less effective treatments, enhance overall welfare, and maintain statistical validity for analyzing experimental data.
-Nevertheless, because CARA designs rely on participant responses collected during the experiment to adjust treatment assignments, delays in observing these responses can compromise the design’s effectiveness.
-This can be intuitively understood as follows: during interim analyses, if the primary outcomes for some participants are not yet observed, the experimental designer may lack the necessary information to optimize treatment assignments for future participants who have profiles similar to those with delayed responses.
-Furthermore, as is often the case in randomized experiments when participants are sequentially enrolled, the response delay mechanisms can vary among participants depending on their assigned treatment arm and baseline covariates.
-Below, we provide a motivating example with arm- and covariate-dependent response delays in a randomized experiment:
-Example 1 (Arm- and covariate-dependent randomized experiments for a cash incentive trial among people living with HIV (PLHIV)).
-In a randomized experiment conducted in Shinyanga region, Tanzania (Fahey et al., 2020), the goal is to assess the impact of cash incentives on improving retention in care and thus achieving viral suppression for PLHIV.
-In this experiment, 530 PLHIV aged 18 years or older were sequentially enrolled and randomly assigned to either receive cash incentives (treatment arm) or no cash incentives (control arm) between April 24, 2018, and December 14, 2018.
-Retention in care is evaluated by individual adherence to antiretroviral therapy measured by viral load six months after treatment assignment.
-However, delays in measuring viral load are observed, as the lab test requires clinical visits for blood draws, and some PLHIV miss their scheduled appointments.
-To investigate these response delay mechanisms, we record the date at which each patient completed six months of treatment and calculate the number of days delayed as the difference between the date of the viral load test and the six-month completion date.
-We further include biological sex as a baseline covariate to examine if the delay mechanism varies across females and males.
-Figure 1: Distribution of the number of delayed days in viral load response collection across treatment arms and baseline covariate (i.e., biological sex) after completing six months of antiretroviral therapy.
-Figure 1 presents a histogram showing the number of delayed days for viral load measurements, stratified by treatment and control arms as well as biological sex.
-The figure illustrates that the distribution of delayed days is influenced by both treatment assignment and biological sex.
-Specifically, the treatment arm shows a higher percentage of responses collected within 100 days compared to the control arm.
-Moreover, within the treatment arm, a larger proportion of female participants have response delays of fewer than 100 days compared to males.
-This observation suggests that for females receiving cash transfers, most viral load responses were collected within three months of completing the treatment.
-Overall, Figure 1 highlights that response delays in this randomized experiment are dependent on both the treatment arm and covariate.
-To our knowledge, although there has been discussion about how response delays might affect statistical inference at the conclusion of CARA(Bai et al., 2008; Hu et al., 2008; Robertson et al., 2023), no previous CARA design has leveraged the delay mechanism to explicitly optimize experimental objectives.
-In this manuscript, we address the challenge of response delays in CARA and seek to answer the following questions: Does the optimal treatment allocation proposed in existing CARA literature remain effective for achieving design objectives?
-If not, what is the optimal treatment allocation?
-How can CARA be adapted to incorporate response delays for more effective attainment of these objectives?
-Finally, at the conclusion of the experiment, how can we construct an efficient statistical estimator that accounts for these delays?
-The following paragraphs outline our contributions in addressing these questions within the context of existing CARA literature.
+1.1 Motivation and contribution
+In clinical trials, social science field experiments, and A/B tests, participants often enroll sequentially, and responses to treatments can vary due to individual-specific characteristics.
+In these contexts, covariate-adjusted response-adaptive (CARA) designs have gained increasing popularity: by dynamically adjusting treatment assignments based on accumulated covariates and outcome information during the experiment, CARA designs can be adopted to optimize treatment allocation to improve statistical power (Hahn, Hirano and Karlan, 2011; Hu, Zhu and Hu, 2015; Blackwell, Pashley and Valentino, 2023), reduce participant exposure to less effective treatments and enhance overall welfare (Hu and Rosenberger, 2006; Hu, Zhu and Hu, 2015; Wei, Ma and Wang, 2024, 2025), and maintain statistical validity for analyzing experimental data (Zhang, Hu, Cheung and Chan, 2007; Hu and Hu, 2012; Bugni, Canay and Shaikh, 2018; Offer-Westort, Coppock and Green, 2021; Bai, Romano and Shaikh, 2022; Robertson, Lee, López-Kolkovska and Villar, 2023; Zhao, 2023; Wei, Ma and Wang, 2025; Bibaut and Kallus, 2025).
+Nevertheless, because CARA designs adjust treatment assignments based on observed participant responses, delays in outcome observation can significantly hinder their effectiveness.
+Intuitively, if primary outcomes for some participants remain unobserved at the time of interim analysis, the experimental designer may lack information needed to optimize treatment assignments for future participants with similar covariate profiles.
+These delays often depend on both the treatment arm and participant covariates, as is common in sequentially enrolled randomized experiments.
+As an example, we revisit the study by Fahey, Njau, Katabaro, Mfaume, Ulenga, Mwenda et al.
+(2020), a randomized experiment conducted in the Shinyanga region of Tanzania between April 24 and December 14, 2018.
+The study aimed to assess the impact of cash incentives on retention in care and viral suppression among people living with HIV (PLHIV).
+A total of 530 PLHIV aged 18 or older were sequentially enrolled and randomly assigned to either a treatment group receiving cash incentives or a control group without any incentive.
+Viral load was measured six months after treatment assignment through blood draws.
+However, delays occurred, as the lab test required a separate clinical visit and some participants missed their scheduled appointments.
+We calculated the number of days delayed as the difference between the date of the viral load test and the six-month follow-up date, separately for two covariate strata defined by biological sex.
+From Figure 1, it is clear that the distribution of outcome delays varies significantly depending on treatment assignment and gender.
+Specifically, the treatment arm exhibits less delay, and within the treatment arm, the delayed response issue seems less pronounced among female participants.
+Overall, the empirical observation highlights the importance of accounting for delayed outcomes that are both arm- and covariate-specific.
+Figure 1: Delays in viral load blood test across treatment arms and covariate strata (biological sex).
+To our knowledge, although there has been prior work on developing valid statistical inference procedures in the presence of delayed responses in CARA (Bai, Hu and Rosenberger, 2002; Hu, Zhang, Cheung and Chan, 2008; Robertson, Lee, López-Kolkovska and Villar, 2023), no CARA designs has incorporated the delay mechanism when optimizing experimental objectives.
+In this paper, we address the challenges that response delays pose to covariate-adjusted response-adaptive (CARA) designs.
+We begin by examining whether existing optimal allocation strategies remain valid under delayed outcomes and, when they do not, derive the optimal allocation in such settings.
+We then develop methods to adapt CARA designs accordingly.
+Finally, we propose an efficient estimator that accounts for delayed responses at the end of the experiment.
+We elaborate on these contributions in more detail below.
 First, we demonstrate that the classical optimal allocation strategy used in existing CARA designs no longer provides optimal treatment assignment when accounting for arm- and covariate-dependent delays.
-Specifically, we first derive a semiparametric efficiency bound for treatment effect estimation accommodating both multi-valued treatments and outcome delay, which differ substantially from existing results in the literature (Hahn, 1998; Hirano et al., 2003; Cattaneo, 2010; Zhu and Zhu, 2023) (see Newey 1990; Van der Vaart 2000; Tsiatis 2006 for general discussions on semiparametric efficiency).
-By deriving semiparametric efficiency bounds that incorporate response delays in multi-stage CARA designs (Theorem 1), we show that the optimal delay mechanism involves a unique trajectory of treatment assignment probabilities that vary across different stages, resulting in treatment allocations that reduce variance for power maximization and minimize failure rates in welfare improvement designs (Figure 2).
-To our knowledge, this is the first design to explicitly explore delay mechanisms to optimize trial objectives, while existing literature focuses on conducting sequential tests (Hampson and Jennison, 2013), designing Bayesian adaptive designs where delay is independent of the arm and covariates (Lin et al., 2020), or conducting statistical inference after experiments with delayed responses have concluded (Bai et al., 2008; Hu et al., 2008).
-Second, because the delay mechanism is unknown a priori and only partially learnable during the experiment, we propose a fully forward-looking CARA design that sequentially updates treatment assignment probabilities to optimize the trial outcome by the experiment’s end while also estimating future delays to incorporate the experimental designer’s preferences.
-Our design encounters two design objectives frequently adopted in current CARA literature: maximizing experiment power (Tymofyeyev et al., 2007; Zhao, 2023) and improving overall participant welfare (or failure reduction) while maintaining statistical power constraints (Rosenberger et al., 2001; Robertson et al., 2023).
+Specifically, we first derive a semiparametric efficiency bound for treatment effect estimation in the presence of outcome delay (Theorem 1), which differ substantially from existing results in the literature (Hahn, 1998; Hirano, Imbens and Ridder, 2003; Cattaneo, 2010; Zhu and Zhu, 2023; Wu, Zheng, Zhang, Zhang and Wang, 2025).
+See Van der Vaart (2000) and Tsiatis (2006) for general discussions on semiparametric efficiency.
+This result allows us to show that the optimal delay mechanism involves a unique trajectory of treatment assignment probabilities that vary across different stages, resulting in treatment allocations with lower variance for power maximization, and enhanced failure reduction in welfare improvement designs.
+To our knowledge, this is the first design to explicitly explore delay mechanisms to optimize trial objectives, while existing literature focuses on conducting sequential tests (Hampson and Jennison, 2013), designing Bayesian adaptive designs where delay is independent of the arm and covariates (Lin, Thall and Yuan, 2020), or conducting statistical inference after experiments have concluded (Bai, Hu and Rosenberger, 2002; Hu, Zhang, Cheung and Chan, 2008).
+Second, since the delay mechanism is unknown a priori and remains only partially estimable during the experiment, we propose a fully forward-looking CARA design that sequentially updates treatment assignment probabilities to target a trial objective at the conclusion of the experiment.
+In this process, our algorithm offers several extrapolation strategies when estimating delay mechanism, allowing better alignment with the designer’s prior knowledge.
+We further discuss in detail two design objectives frequently adopted in current CARA literature: maximizing experiment power (Tymofyeyev, Rosenberger and Hu, 2007; Zhao, 2023) and improving overall participant welfare (or failure reduction) while maintaining statistical power constraints (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001; Robertson, Lee, López-Kolkovska and Villar, 2023).
 We justify the benefits of our proposed forward-looking CARA design through theoretical investigation (Section 4) and a synthetic case study (Section 5).
 Under the design objective of maximizing experiment power, our proposed forward-looking CARA design attains higher estimation efficiency compared to the Neyman allocation and complete randomization designs in the presence of delayed responses.
-The efficiency gain is further justified by Theorem 5.
-Under the design objective of improving overall participant welfare, we demonstrate that our proposed forward-looking CARA design achieves a higher failure reduction rate, leading to greater welfare improvement compared to both the ethical design and the complete randomization design while maintaining statistical power.
-For both design objectives, our proposed design can provide valid statistical inference, which is justified in Theorem 2 and Theorem 3 and further demonstrated in a synthetic case study (Section 5).
-Third, we make theoretical contributions to the literature on CARA and adaptive experimental designs by establishing general conditions under which estimated treatment effects are consistent and asymptotically normally distributed (Lemma 1 and Theorem 2).
+Under the design objective of improving overall participant welfare, we demonstrate that our proposed forward-looking CARA design achieves additional failure reduction, leading to greater welfare improvement compared to both the ethical design and the complete randomization design while maintaining statistical power.
+For both design objectives, our proposed design can provide valid statistical inference, which is justified in Theorem 3.
+Third, we make theoretical contributions to the literature on CARA and adaptive experimental designs by establishing general conditions under which estimated treatment effects are consistent and asymptotically normally distributed (Theorem 2).
 Specifically, we do not restrict the potential outcomes or the delay mechanism to any parametric family of distributions, thereby alleviating the burden of choosing a particular set of parametric assumptions.
-Such flexibility is crucial for many applications, as the delay mechanism is typically arm- and covariate-dependent, as our earlier examples have demonstrated.
-Our theoretical development builds on martingale methods (Hall and Heyde, 1980) and a high-level condition (Condition 1) that the optimized allocations converge in large samples—an assumption satisfied by many adaptive experimental designs.
+Our theoretical development builds on martingale methods (Hall and Heyde, 1980) and a high-level condition (Condition 1) that the optimized allocations converge in large samples, an assumption satisfied by many adaptive experimental designs.
 As a result, the general consistency and asymptotic normality findings may be of independent interest.
 We then specialize in our proposed design and verify the design consistency condition (Section 4.3), providing a suite of tools for designing and analyzing adaptive experiments with delayed outcomes, along with rigorous statistical guarantees.
-1.2 Setup: CARA with arm and covariate dependent delayed outcomes
+1.2 Setup: CARA with arm- and covariate-dependent delays
 In the remainder of this introduction, we lay out the setup of CARA when responses are subject to delay.
@@
 Participants are indexed by $i=1,2,\dots,N$, with $N=\sum_{t=1}^{T}n_{t}$ being the total sample size.
-Occasionally, we will use $N_{t}$ to denote the cumulative sample size at the end of stage $t$, that is, $N_{t}=\sum_{s=1}^{t}n_{s}$.
+We use $N_{t}$ to denote the cumulative sample size at the end of stage $t$, that is, $N_{t}=\sum_{s=1}^{t}n_{s}$.
 Therefore, $N=N_{T}$, where we omit the subscript to save notation.
 Finally, $S_{i}=t$ denotes that individual $i$ is enrolled in stage $t$.
-Upon enrollment, baseline covariates $X_{i}\in\mathcal{X}$ are collected for each participant, and they are then randomized into one of the treatment arms, denoted by $A_{i}\in\mathcal{A}$, which represents their actual treatment status.
-Following the Neyman-Rubin causal model, the potential outcomes are denoted by $Y_{i}(a)$ for different treatment $a\in\mathcal{A}$.
-In line with existing literature in adaptive design, we are in a scenario that the potential outcomes and the covariates, $(Y_{i}(a),X_{i})$, are independently and identically distributed across different stages.
-For future reference, we also introduce notation for mean potential outcomes: $\mu(x,a)=\mathbb{E}[Y_{i}(a)|X_{i}=x]$ and $\mu(a)=\mathbb{E}[Y_{i}(a)]$, [^1] Other moments of the potential outcomes are $\mu_{s}(x,a)=\mathbb{E}[Y_{i}(a)^{s}|X_{i}=x]$ and $\mu_{s}(a)=\mathbb{E}[Y_{i}(a)^{s}]$, but we omit the subscript whenever $s=1$ for ease of notation.
-and the conditional and unconditional treatment effects are defined as
-$\tau(x,a^{\prime},a)$ $=\mu(x,a^{\prime})-\mu(x,a)\quad\text{and}\quad\tau(a^{\prime},a)=\mu(a^{\prime})-\mu(a).$
+Upon enrollment, baseline covariates $X_{i}\in\mathcal{X}$ are collected for each participant, and they are then randomized into one of the treatment arms, denoted by $A_{i}\in\mathcal{A}=\{0,1\}$, which represents their actual treatment status.
+Following the Neyman-Rubin causal model, the potential outcomes are denoted by $Y_{i}(0)$ and $Y_{i}(1)$.
+In line with existing literature in adaptive design, we are in a scenario where the potential outcomes and the covariates, $(Y_{i}(0),Y_{i}(1),X_{i})$, are independent and identically distributed across different stages.
+For future reference, we also introduce notation for mean potential outcomes: $\mu(x,a)=\mathbb{E}[Y_{i}(a)|X_{i}=x]$ and $\mu(a)=\mathbb{E}[Y_{i}(a)]$, and the conditional and unconditional treatment effects are defined as $\tau(x)=\mu(x,1)-\mu(x,0)$ and $\tau=\mu(1)-\mu(0)$.
+Other moments of the potential outcomes are $\mu_{s}(x,a)=\mathbb{E}[Y_{i}(a)^{s}|X_{i}=x]$ and $\mu_{s}(a)=\mathbb{E}[Y_{i}(a)^{s}]$, but we omit the subscript whenever $s=1$ for ease of notation.
 Since outcome information may not be immediately available after treatment assignment, we use $D_{i}$ to represent the number of stages after which the experimenter can observe $Y_{i}$.
@@
 Also note that when $D_{i}=0$, it implies no delay in the outcome for this individual.
-In this manuscript, we allow the delay mechanism $D_{i}$ to depend on the covariates and the treatment assignment status, and is independent of the potential outcomes conditional on the covariates $X_{i}$ and the treatment assignment $A_{i}$.
-Similar to the assumption of identical distribution of potential outcomes across stages mentioned earlier, we assume that the delay distribution does not shift over time.
-Together, these can be formally expressed as:
 […]
-As the delay distribution does not shift over time, it allows us to sequentially learn the delay mechanism using accrued data and to optimize treatment allocation.
-Finally, we note that this assumption does not preclude the possibility that the delay mechanism may depend on the observed outcomes.
 In CARA designs, the treatment assignment is sequentially updated based on accumulated data to achieve a pre-specified design objective.
-Specifically, at stage $t$, suppose we have collected historical participant information, denoted as
-$\mathcal{H}_{t}=\{(A_{i},X_{i},Y_{i},D_{i})\}_{i=1}^{N_{t}},$
-where we recall that the sample indexed by $i=1,2,\dots,N_{t}$ consists of all participants enrolled in the first $t$ stages.
+Specifically, at stage $t$, suppose we have collected historical participant information, denoted as $\mathcal{H}_{t}=\{(A_{i},X_{i},Y_{i},D_{i})\}_{i=1}^{N_{t}}$, where we recall that the sample indexed by $i=1,2,\dots,N_{t}$ consists of all participants enrolled in the first $t$ stages.
 The treatment assignment for the next stage, $t+1$, denoted as $\hat{e}^{*}_{t+1}(a|x)$, is determined by both this historical information and the covariate information of the newly enrolled participants in Stage $t+1$.
-Formally, this is captured by
-$\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t+1,\mathcal{H}_{t}]=\hat{e}^{*}_{t+1}(a|x),$
-which means the treatment assignment adapts to $\mathcal{H}_{t}$.
+Formally, this is captured by $\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t+1,\mathcal{H}_{t}]=\hat{e}^{*}_{t+1}(a|x)$, which means the treatment assignment adapts to $\mathcal{H}_{t}$.
+In this manuscript, we work under the setting that $T$ and $n_{t}$ are pre-fixed.
+We also hope to note that, in general adaptive trial designs, determining the appropriate sample size and the number of stages involves balancing statistical rigor, operational feasibility, and ethical considerations.
+Before an experiment starts, practitioners often pre-specify the desired power and significance level, along with the expected treatment effect size based on prior studies or clinical knowledge, to compute an initial overall sample size.
+Then, the number of stages is selected by evaluating logistical considerations (e.g., recruitment rate, data availability, cost constraints).
+It is often the case that fewer stages are used to simplify logistics and reduce operational complexity, while more stages increase flexibility but impose higher management costs.
+Once the number of stages is decided, sample size allocation per stage can be determined either equally or unequally (e.g., adaptive design with small initial pilot study) based on anticipated interim analyses.
+It is also common practice to conduct simulations during the planning stage to evaluate various stage/sample size configurations, ensuring feasibility and ethical appropriateness of the chosen adaptive strategy.
+Our design thus aligns with the case where there is a finite number of stages and each stage contains an equal/unequal number of participants.
 2 Optimal delay-adjusted treatment allocation in CARA
-In this section, we discuss how delayed response in CARA affects optimal treatment allocation rules.
-Compared to the existing literature on CARA or response adaptive design (RAR) with delayed outcomes (Bai et al., 2008; Hu et al., 2008), to our knowledge, no prior designs have explicitly leveraged delay mechanisms to optimize experimental objectives.
-While the design goals of CARA vary across applications, maintaining statistical power to assess the treatment’s effectiveness on the primary outcome remains a central concern.
-A statistically efficient estimator of the treatment effect also contributes to a robust adaptive design, as the estimated treatment effects are used to sequentially revise the treatment allocation strategy.
-However, it remains unclear what constitutes a good estimator of the treatment effect in the presence of delayed outcomes.
-Moreover, the semiparametric efficiency bound for estimating the treatment effect in multi-stage experiments with delayed outcomes has not been developed.
-To provide practical guidance for designing response-adaptive experiments, we first present a new result on the semiparametric efficiency bound of estimating the average treatment effect, defined as the mean difference between the potential outcomes of two treatment arms $a$ and $a^{\prime}$.
-In such experiments with delayed outcomes, we have
-$\int\limits_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,a^{\prime})}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a^{\prime}|x)\rho(T-t|x,a^{\prime})}$ $+\frac{\sigma^{2}(x,a)}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}+\big{(}\tau(x,a^{\prime},a)-\tau(a^{\prime},a)\big{)}^{2}\Bigg{]}\mbox{d}x,$ (1)
-where $r_{t}=\mathbb{P}[S_{i}=t]$, the fraction of participants enrolled in stage $t$, $p(x)$ is the probability density function of the covariate $X$, and $\sigma_{t}^{2}(x,a)$ is the conditional variance of potential outcomes under treatment $a\in\{0,1\}$ at time $t$.
-See Section 4, Theorem 1 for a rigorous statement.
-From the above result, it is evident that when the delay mechanism of the response depends on the covariates and treatment status, the optimal treatment allocation rules in classical CARA designs must be modified to account for the impact of the delay mechanism.
-In what follows, we examine how the optimal treatment allocation rule is affected by the arm- and/or covariate-dependent delayed responses in two common experimental objectives: one aimed at enhancing statistical power (Tymofyeyev et al., 2007), and the other focused on minimizing the expected number of failures while maintaining a minimum power requirement for binary outcomes (Rosenberger et al., 2001).
+While the design goals of CARA vary across applications, maintaining sufficient statistical power to assess the treatment’s effectiveness on the primary outcome remains a central concern.
+A statistically efficient estimator of the treatment effect also supports a robust adaptive design, as the estimated effects guide the sequential revision of the treatment allocation strategy.
+However, it remains unclear what constitutes a good estimator in the presence of delayed outcomes.
+Moreover, the semiparametric efficiency bound for estimating the treatment effect in multi-stage experiments with delayed responses has not yet been established.
+To provide practical guidance for designing response-adaptive experiments, we begin by presenting a new result on the semiparametric efficiency bound for estimating the average treatment effect:
+$\int\limits_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,1)}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|x,1)e_{t}(1|x)}$ $+\frac{\sigma^{2}(x,0)}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|x,0)e_{t}(0|x)}+\big{(}\tau(x)-\tau\big{)}^{2}\Bigg{]}\mbox{d}x,$ (1)
+where $r_{t}=\mathbb{P}[S_{i}=t]$ being the fraction of participants enrolled in stage $t$, $p(x)$ is the probability density function of the covariate $X$, and $\sigma^{2}(x,a)$ is the conditional variance of the potential outcome $Y_{i}(a)$.
+See Theorem 1 in Section 4 below for a rigorous statement.
+To intuitively understand Eq (1), it helps to first revisit the classical semiparametric efficiency bound in a static (one-period) setting without delays, which is given by:
+$\int\limits_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,1)}{e(1|x)}$ $+\frac{\sigma^{2}(x,0)}{e(0|x)}+\big{(}\tau(x)-\tau\big{)}^{2}\Bigg{]}\mbox{d}x.$ (2)
+In this classical formulation, the denominators (i.e, propensity scores $e_{t}(a|x)$) adjust for covariate-specific treatment assignment.
+The key difference between Eq (2) and Eq (1) is the introduction of an additional term $\rho(T-t|x,a)$ in the denominators, which accounts for the probability distribution of response delays.
+Intuitively, delays introduce extra uncertainty or missingness in observing outcomes.
+As we operate in a multiple-stage experimental setting, we aggregate these delay-related probabilities across all stages, appropriately weighted by each stage’s sample size captured by $r_{t}$.
+Hence, the revised efficiency bound appropriately adjusts for both the random assignment of treatments (captured by propensity scores) and the randomness arising from delays in observing responses.
+From the above result, it is evident that when the delay mechanism of the response depends on the covariates and treatment status, the optimal treatment allocation rules in classical CARA designs must be modified to account for outcome delays.
+In what follows, we examine how the optimal treatment allocation rule is affected by the arm- and/or covariate-dependent delayed responses in two common experimental objectives: one aimed at enhancing statistical power (Tymofyeyev, Rosenberger and Hu, 2007), and the other focused on minimizing the expected number of failures while maintaining a minimum power requirement for binary outcomes (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001).
 In both design goals, the variance of the estimated treatment effects plays an essential role.
-To simplify the presentation, we focus on two-arm CARA designs with $a\in\{0,1\}$.
-For multi-arm experiments, the design objective can be generalized using noncentrality parameters, as discussed in Lachin (1977); Tymofyeyev et al.
-(2007).
-Design objective 1: Optimal treatment allocation in power maximization design with delayed response.
-This design aims to sequentially assign treatments to minimize the variance of the estimated average treatment effect (ATE) for two treatment arms, denoted by $a\in\{0,1\}$, thereby maximizing statistical power.
-The optimal allocation is determined by solving the following optimization problem:
-$\min_{e_{1},\ldots,e_{T}}\ \mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))\quad\text{subject to}\quad\delta\leq e_{t}(\cdot)\leq 1-\delta\quad\text{for }t=1,\ldots,T,$ (2)
-where $\delta$ ensures that the assignment probabilities are bounded away from 0 and 1, maintaining sufficient randomness in treatment assignments.
+To simplify the presentation, we focus on two-arm CARA designs with $a\in\{0,1\}$ in this paper.
+For multi-arm experiments, the design objective can be generalized using non-centrality parameters, as discussed in Lachin (1977) and Tymofyeyev, Rosenberger and Hu (2007).
+Design objective 1: Optimal treatment allocation for power maximization.
+The objective is to sequentially assign treatments to minimize the variance of the estimated average treatment effect (ATE), thereby maximizing statistical power.
+The optimal allocation is the solution to the optimization problem:
+$\min_{e_{1}(\cdot),\dots,e_{T}(\cdot)}\ \mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))\quad\text{subject to}\quad\delta\leq e_{t}(\cdot)\leq 1-\delta\quad\text{for }t=1,\ldots,T,$ (3)
+where $\delta$ ensures that the assignment probabilities are bounded away from 0 and 1, maintaining sufficient randomness in treatment assignments (Ma and Wang, 2020; Heiler and Kazak, 2021; Sasaki and Ura, 2022; Ma, Sasaki and Wang, 2025; Dorn, 2025).
 The objective function $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ is a component of the semiparametric efficiency bound related to the propensity scores, and is defined as:
-$\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))=\sum_{x\in\mathcal{X}}p(x)\left[\frac{\sigma_{t}^{2}(x,1)}{\sum_{t=1}^{T}r_{t}e_{t}(1|x)\rho(T-t|x,1)}+\frac{\sigma_{t}^{2}(x,0)}{\sum_{t=1}^{T}r_{t}e_{t}(0|x)\rho(T-t|x,0)}\right].$
+$\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))=\sum_{x\in\mathcal{X}}p(x)\left[\frac{\sigma_{t}^{2}(x,1)}{\sum_{t=1}^{T}r_{t}\rho(T-t|x,1)e_{t}(x)}+\frac{\sigma_{t}^{2}(x,0)}{\sum_{t=1}^{T}r_{t}\rho(T-t|x,0)(1-e_{t}(x))}\right].$
 In line with existing literature on CARA designs, we explicitly consider discrete covariates in the objective function, as most design applications will first group participants based on their covariate information.
 Therefore, the objective function takes a summation form with respect to the probability mass function.
-Design objective 2.
-Optimal treatment allocation in design on failure reduction under a power constraint.
-This design aims to minimize the expected number of failures while satisfying a minimum power requirement for binary responses (Rosenberger et al., 2001).
+Design objective 2: Optimal treatment allocation for failure reduction subject to a power constraint.
+This design aims to minimize the expected number of failures while satisfying a minimum power requirement for binary responses (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001).
 In this context, a treatment is considered to fail when the potential outcome equals one.
 The optimal allocation thus seeks to assign treatments based on the solution to the following optimization problem:
-$\min_{e_{1},\ldots,e_{T}}$ $\ \sum_{t=1}^{T}r_{t}\sum_{x\in\mathcal{X}}p(x)\Big{[}e_{t}(1|x)\left(1-\mu(x,1)\right)+\left(1-e_{t}(0|x)\right)\left(1-\mu(x,0)\right)\Big{]}$ (3)
-s.t. $\ \delta\leq e_{t}(\cdot)\leq 1-\delta,\quad\text{for }t=1,\ldots,T$
-$\ \mathsf{V}\left(e_{1}(\cdot),\dots,e_{T}(\cdot)\right)+\sum_{x\in\mathcal{X}}p(x)\big{(}\tau(x,1,0)-\tau(1,0)\big{)}^{2}\leq C$
+$\min_{e_{1}(\cdot),\dots,e_{T}(\cdot)}\ \sum_{t=1}^{T}r_{t}\sum_{x\in\mathcal{X}}p(x)\Big{[}e_{t}(x)\left(1-\mu(x,1)\right)+\left(1-e_{t}(x)\right)\left(1-\mu(x,0)\right)\Big{]}\quad$ (4)
+$\qquad\qquad\text{s.t.}\quad\ \delta\leq e_{t}(\cdot)\leq 1-\delta,\quad\ \mathsf{V}\left(e_{1}(\cdot),\dots,e_{T}(\cdot)\right)+\sum_{x\in\mathcal{X}}p(x)\big{(}\tau(x)-\tau\big{)}^{2}\leq C.$
 In this formulation, the objective function minimizes the total expected number of failures during the entire experiment.
-The term inside the summation accounts for the expected failures when assigning treatment 1 with probability $e_{t}(\cdot)$ and treatment 00 with probability $1-e_{t}(\cdot)$.
+The term inside the summation accounts for the expected failures when assigning treatment 1 with probability $e_{t}(x)$ and treatment 0 with probability $1-e_{t}(x)$.
 Here, $\mu(x,a)$ represent the expected success probabilities under treatments $a$, given covariates $x$.
-The first feasibility constraints again ensure that the treatment assignment probabilities remain within the interval $[\delta,1-\delta]$.
+The first feasibility constraints again ensure that the treatment assignment probabilities remain within the range $\delta,1-\delta]$.
 This design balances the ethical imperative to reduce adverse outcomes with statistical consideration to maintain sufficient power for detecting treatment effects.
 Due to the delay mechanism’s influence on the variance lower bound, the optimal treatment allocation rules differ substantially from those derived in the existing literature.
-We shall show that the solutions to the optimization problems in Eq (2) and Eq (3), denoted as $e_{1}^{*}(\cdot),\ldots,e_{T}^{*}(\cdot)$, from a sequence of probabilities that vary across different stages.
+We shall show that the solutions to the optimization problems in (3) and (4), denoted as $e_{1}^{*}(\cdot),\ldots,e_{T}^{*}(\cdot)$, form a sequence of probabilities that vary across different stages.
 In comparison, the classical Neyman allocation rule for power maximization:
-${e}_{t}^{\mathtt{Neyman}}(a|x)=\frac{\sigma(x,a)}{\sigma(x,1)+\sigma(x,0)},\quad a\in\{0,1\},\quad t=1,\ldots,T,$ (4)
-and the optimal failure reduction rule derived in Rosenberger et al. (2001):
-${e}_{t}^{\mathtt{FR}}(a|x)=\frac{\sqrt{\mu(x,a)}}{\sqrt{\mu(x,1)}+\sqrt{\mu(x,0)}},\quad a\in\{0,1\},\quad t=1,\ldots,T,$ (5)
+$e_{t}^{\mathtt{Neyman}}(a|x)=\frac{\sigma(x,a)}{\sigma(x,1)+\sigma(x,0)},\quad\text{for }t=1,\ldots,T,$ (5)
+and the optimal failure reduction rule of Rosenberger, Stallard, Ivanova, Harper and Ricks (2001):
+$e_{t}^{\mathtt{FR}}(a|x)=\frac{\sqrt{\mu(x,a)}}{\sqrt{\mu(x,1)}+\sqrt{\mu(x,0)}},\quad\text{for }t=1,\ldots,T,$ (6)
 are both fixed probabilities that do not differ across stages.
@@
 Therefore, when response delay plays a significant role, assigning treatments based on existing optimal rules no longer guarantees achieving the desired experimental goals.
-Figure 2: Comparison between classical optimal rule and delay-adjusted optimal rule for two design objectives in CARA: (A)-(B) power maximization and (C)-(D) reducing the total number of failures.
-We illustrate this point more concretely in a simulation study, where model parameters are calibrated from our synthetic case study and are applied to the oracle optimization problems in Eq (2) and (3).
-We refer to the detailed data-generating process for this simulation design in Section 5.
-First, we solve for the oracle treatment allocation with the design objective of maximizing power, as defined in Eq (2).
-The resulting oracle treatment allocations for each Stage $t$ and the Neyman allocation are presented in Panels (A) and (B) for covariate strata $X=0$ and $X=1$, respectively.
-The corresponding values of the objective function in Eq (2) under the two different allocations are labeled as $\mathsf{V}_{\text{Delay}}$ and $\mathsf{V}_{\text{Neyman}}$.
-Second, we solve for the oracle treatment allocation with the design objective of minimizing the total expected number of failures, as in Eq (3).
-The oracle treatment allocation and the ethical allocation across Stage $t$ are shown in Panels (C) and (D) for the covariate strata $X=0$ and $X=1$, respectively.
-The percentage of failure under the two allocation strategies is labeled in the figures.
-As illustrated in Figure 2(A), the optimal treatment allocation accounting for the response delay (i.e., the solution to Eq (2)) varies across stages and deviates significantly from Neyman allocation.
-Furthermore, the optimal treatment allocations result in an ATE estimator with reduced variance (e.g., $\mathsf{V}_{\text{Delay}}=1.62$ and $\mathsf{V}_{\text{Neyman}}=2.95$ for the strata with $X=0$).
-As illustrated in Figure 2(B), a similar observation applies to ethical designs, where the optimal treatment allocation tailored to the ethical design objective (i.e., the solution to Eq (3)) differs from the ethical design rule in Eq (5), and the delay-adjusted optimal allocation achieves a lower failure rate.
-Before concluding this section, we note that under appropriate conditions, we will show that the optimal treatment allocation rules $e_{1}^{*}(\cdot),\ldots,e_{T}^{*}(\cdot)$ are unique (Theorem 4).
-This uniqueness lays the foundation for establishing the statistical guarantees of our proposed design.
-Moreover, it demonstrates the necessity of accounting for delay mechanisms to achieve the desired experimental design goals.
+Figure 2: Optimal oracle allocation (accounting for outcome delays) and the Neyman allocation for the power maximization objective.
+We illustrate this point more concretely by computing the optimal oracle treatment allocation in the example introduced in Section 1, and compare it with classical treatment assignment methods such as the Neyman allocation.
+The covariate strata we consider are formed by two covariates, the biological sex and the WHO clinical stages of HIV.
+See Section 5 for additional details on the model parameters.
+In Figure 2, we present the oracle allocations for each stage and each covariate stratum corresponding to the power maximization objective defined in (3), as well as the Neyman allocation.
+Note that, since the classical Neyman allocation ignores the delayed outcome issue, it results in a constant treatment probability across the experimental stages, whereas the oracle allocation varies significantly across both stages and covariate strata.
+This highlights the importance of allowing the delay mechanism to be both arm- and covariate-specific.
+We also report the objective function value for each design method in the figure.
+As can be seen, the efficiency gain in the oracle allocation can be substantial: more than 10% in the first covariate stratum and about 8% in the last case.
+Once again, this numerical exercise demonstrates the potential gains from leveraging the delay mechanism in a CARA framework.
 3 Fully forward-looking CARA with delayed outcomes
@@
 In practice, the delay mechanism is rarely known and must be estimated as data are sequentially accumulated to optimize the experimental objective.
-In this section, we propose two delay-informed practical CARA designs tailored to two previously discussed experimental goals: improving statistical power (Tymofyeyev et al., 2007) and minimizing the expected number of failures while meeting a minimum power requirement for binary responses (Rosenberger et al., 2001).
+In this section, we propose two delay-informed practical CARA designs tailored to two previously discussed experimental goals: improving statistical power (Tymofyeyev, Rosenberger and Hu, 2007) and minimizing the expected number of failures while meeting a minimum power requirement for binary responses (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001).
 We also provide tools for valid statistical inference after the experiment is finished.
-As with any CARA designs, at the first stage of the experiment, we lack information about the participants, so treatments are assigned randomly with a pre-fixed probability.
-In subsequent stages, our proposed fully forward-looking CARA design incorporates two key components: (1) estimating past delay mechanisms from collected data while anticipating future delays, and (2) using these estimates to formulate a fully forward-looking optimization problem that determines optimal treatment assignments for the next stage.
-An overview of our proposed fully forward-looking CARA can be found in Figure 3.
-We introduce these two components in the following paragraphs.
-To save notation, we consider a slightly simplified setting in this section that there are only two arms, that is, $a\in\mathcal{A}=\{0,1\}$.
-We remark that our design strategies can be easily extended to a more general multi-valued treatment setting, with the cost of additional notation.
-In Section 4, we will establish properties of our proposed design methods and the treatment effect estimators under the general setting, accommodating multiple treatment arms.
-Figure 3: Forward looking CARA design illustration.
-At the end of each stage $t\geq 2$, we estimate (and update) the observed delay probabilities as
-$\hat{\rho}_{t}(d|x,a)=\sum_{\ell=0}^{d}\hat{\mathbb{P}}[D_{i}=\ell|X_{i}=x,A_{i}=a]=\sum_{\ell=0}^{d}\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}=\ell,S_{i}\leq t-\ell)}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,S_{i}\leq t-\ell)}},\quad d\leq t-1,$
+As with any CARA design, we use a pre-specified and fixed treatment assignment rule for the first stage due to the lack of a priori knowledge about the covariate and outcome distributions as well as the delay mechanism.
+In subsequent stages, our proposed fully forward-looking CARA design incorporates two key components: (i) estimating aspects of the delay mechanism from accrued data, while employing extrapolation to assess the tail behavior of the delay distribution; and (ii) using these estimates to formulate a feasible, fully forward-looking optimization problem that determines the optimal treatment assignments for the next stage.
+We first discuss the estimation of the delay mechanism. Assuming stage $t$ of the experiment has concluded, we first update our delay distribution estimator via
+$\hat{\rho}_{t}(d|x,a)=\sum_{\ell=0}^{d}\hat{\mathbb{P}}[D_{i}=\ell|X_{i}=x,A_{i}=a]=\sum_{\ell=0}^{d}\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}=\ell,S_{i}\leq t-\ell)}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,S_{i}\leq t-\ell)}},\quad d\leq t-1,$ (7)
 for $a=0,1$.
@@
 To gain some intuition, consider a concrete example for $t=3$.
-At the end of the third stage, data collected from participants enrolled in the first stage will be informative about the delay mechanism $\rho(0)$, $\rho(1)$, and $\rho(2)$.
+At the end of the third stage, data collected from participants enrolled in the first stage will be informative about the delay mechanism $\rho(0|\cdot)$, $\rho(1|\cdot)$, and $\rho(2|\cdot)$.
 Among those whose outcome information is still missing at the end of stage 3, it is not possible to tell whether such information will become available until the end of stage 4.
-Following a similar reasoning, participants enrolled in the second stage can help estimate $\rho(0)$ and $\rho(1)$.
+Following a similar reasoning, participants enrolled in the second stage can help estimate $\rho(0|\cdot)$ and $\rho(1|\cdot)$.
 The main message is that at any stage $t$, the delay mechanism is only estimable up to $d\leq t-1$, and the remaining “tail features” are not estimable until more data is collected.
@@
 For a conservative experimenter, we consider a simple extrapolation approach:
-$\hat{\rho}_{t}(d|x,a)=\hat{\rho}_{t}(t-1|x,a),\quad d\geq t.$
+$\hat{\rho}_{t}(d|x,a)=\hat{\rho}_{t}(t-1|x,a),\quad d\geq t.$ (8)
 This extrapolation essentially assumes that all future delay probabilities will remain at the last estimated value.
-For an optimistic experimenter, we consider an extrapolation that postulates all missing outcomes will be available after one more experimental stage:
-$\hat{\rho}_{t}(d|x,a)=1,\quad d\geq t.$
+As we discuss below in Section 4.3, the conservative extrapolation can be understood from a minimax perspective, as we will be sequentially optimizing the objective function under the worst-case scenario of the delay mechanism.
+For this reason, we adopt this approach in most of our numerical experiments and theoretical investigations.
+For an optimistic experimenter, we consider an extrapolation that postulates all missing outcomes will be available after one more experimental stage: $\hat{\rho}_{t}(d|x,a)=1,d\geq t$.
 Lastly, a neutral experimenter might gradually adjust the estimated delay probabilities from the last observed value to immediate response by interpolating between $\hat{\rho}_{t}(t-1|x,a)$ and $1$ for $d=t,\dots,T-1$.
-These estimation strategies enable experimenters to incorporate future delay probabilities into our design, accommodating different perspectives on how delays may evolve.
-An illustration of different delay mechanism estimation strategies is plotted in Figure 4.
+These strategies enable experimenters to incorporate future delay probabilities into our design, accommodating different perspectives on how delays may evolve.
 Although it is generally impossible to rank the three extrapolation methods, we show through Monte Carlo experiments that they all deliver valid inference for the desired causal effects and more effectively achieves the design objectives compared to classical approaches (such as complete randomization).
-To avoid unnecessary length discussions, we focus on the conservative approach in our theoretical investigation and show that the optimized treatment allocation is consistent (Theorem 4).
-Figure 4: Comparison of different experimenter perspectives on delays and their corresponding estimation on the delay mechanism.
 Next, to optimize treatment allocation for stage $t+1$ while accounting for response delays, we introduce a fully forward-looking estimation of the variance term $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ at the current stage $t$.
-Our proposed estimate incorporates past allocations up to Stage $t$, future allocations from stages $t+1$ to $T$, and both past and future delay mechanisms.
-It is designed to anticipate the impact of future allocations and outcome delays on the variance of the final treatment effect estimate after the experiment is finished.
-Our estimate is of the following form:
-$\hat{\mathsf{V}}^{\text{forward}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Bigg{[}\frac{\hat{\sigma}_{t}^{2}(x,1)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)\hat{e}^{*}_{\ell}(1|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)e_{\ell}(x)}$
-$\qquad\quad+\frac{\hat{\sigma}_{t}^{2}(x,0)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)\hat{e}^{*}_{\ell}(0|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)(1-e_{\ell}(x))}\Bigg{]}.$
-In this expression, $\hat{p}_{t}(x)$ represents the estimated proportion of participants with covariate $x$ up to stage $t$, calculated as
-$\hat{p}_{t}(x)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,S_{i}\leq t)}}{N_{t}},$
-where recall that $n_{\ell}$ is the number of participants enrolled at stage $\ell$, $N_{t}$ is the total number of participants up to stage $t$, and $N=N_{T}$.
+Our proposed estimate incorporates past allocations up to Stage $t$, future allocations from stages $t+1$ to $T$, and the estimated delay mechanisms.
+Our method is designed to anticipate the impact of future allocations and outcome delays on the variance of the final treatment effect estimate after the experiment concludes:
+$\hat{\mathsf{V}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Bigg{[}\frac{\hat{\sigma}_{t}^{2}(x,1)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)\hat{e}^{*}_{\ell}(1|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)e_{\ell}(x)}$
+$\qquad\qquad+\frac{\hat{\sigma}_{t}^{2}(x,0)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)\hat{e}^{*}_{\ell}(0|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)(1-e_{\ell}(x))}\Bigg{]}.$
+Here, $\hat{p}_{t}(x)=N_{t}^{-1}\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,S_{i}\leq t)}$, which is the estimated proportion of participants with covariate $x$ up to stage $t$.
 The term $\hat{r}_{\ell}=n_{\ell}/N$ denotes the proportion of participants enrolled at stage $\ell$.
-The estimated conditional variance of the outcome $Y(a)$ given covariate $x$ and treatment $a$, up to stage $t$, is denoted by $\hat{\sigma}_{t}^{2}(x,a)$ and is calculated as
-$\hat{\sigma}_{t}^{2}(x,a)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}(Y_{i}-\hat{\mu}_{t}(x,a))^{2}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}},\quad\hat{\mu}_{t}(x,a)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}Y_{i}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}}.$
-While the expressions may seem complicated, they are simply the sample mean and sample variance of the outcome variable, restricted to the subsample for group $x$ receiving treatment $a$.
+The estimated conditional mean and variance of the outcome $Y_{i}(a)$ are
+$\hat{\sigma}_{t}^{2}(x,a)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}(Y_{i}-\hat{\mu}_{t}(x,a))^{2}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}},\ \hat{\mu}_{t}(x,a)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}Y_{i}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}}.$ (9)
+While the expressions may seem complicated, they are simply the sample mean and sample variance of the outcome, restricted to the subsample for group $x$ receiving treatment $a$.
+We remark that while the true underlying conditional means $\mu(x,a)$ and conditional variances $\sigma^{2}(x,a)$ do not depend on a specific stage $t$, their estimate are stage-specific.
+This is due to the sequential nature of the experiment: we always update previous estimates whenever new data become available.
 Using this fully forward-looking variance estimate, we can then determine an effective treatment assignment for stage $t+1$ that accounts for how future treatment allocations and response delays from stages $t+1$ onward affect the design objectives at the end of the experiment.
 For the first design objective of maximizing power, we solve the following optimization problem to find the treatment allocation probability for stage $t+1$:
-$(\tilde{e}_{t+1}(\cdot),\dots,\tilde{e}_{T}(\cdot))=\operatorname*{argmin}_{e_{t+1},\ldots,e_{T}}\left\{\hat{\mathsf{V}}^{\text{forward}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot)):\ \delta\leq e_{\ell}(\cdot)\leq 1-\delta,\ \ell=t+1,\ldots,T\right\}.$ (6)
-We then set $\hat{e}^{*}_{t+1}(\cdot)=\tilde{e}_{t+1}(\cdot)$.
+$(\tilde{e}_{t+1}(\cdot),\dots,\tilde{e}_{T}(\cdot))=\operatorname*{argmin}_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}\ \hat{\mathsf{V}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot)),$ (10)
+subject to the constraint that $\delta\leq e_{\ell}(\cdot)\leq 1-\delta$ for $\ell\geq t+1$. We then set $\hat{e}^{*}_{t+1}(1|\cdot)=1-\hat{e}^{*}_{t+1}(0|\cdot)=\tilde{e}_{t+1}(\cdot)$.
 For the second design objective of reducing failures under a power constraint, we define a similar fully forward-looking estimate of the expected proportion of total failures as
-$\hat{\mathsf{P}}^{\text{forward}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)=$ $\sum_{\ell=1}^{t}\hat{r}_{\ell}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\left[\hat{e}^{*}_{\ell}(1|x)\left(1-\hat{\mu}_{t}(x,1)\right)+\hat{e}^{*}_{\ell}(0|x)\left(1-\hat{\mu}_{t}(x,0)\right)\right]$
-$+\sum_{\ell=t+1}^{T}r_{l}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\left[e_{t}(1|x)\left(1-\hat{\mu}_{t}(x,1)\right)+e_{t}(0|x)\left(1-\hat{\mu}_{t}(x,0)\right)\right].$
+$\hat{\mathsf{P}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)=$ $\sum_{\ell=1}^{t}\hat{r}_{\ell}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Big{[}\hat{e}^{*}_{\ell}(1|x)\left(1-\hat{\mu}_{t}(x,1)\right)+\hat{e}^{*}_{\ell}(0|x)\left(1-\hat{\mu}_{t}(x,0)\right)\Big{]}$
+$\quad+\sum_{\ell=t+1}^{T}r_{l}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Big{[}e_{\ell}(x)\left(1-\hat{\mu}_{t}(x,1)\right)+(1-e_{\ell}(x))\left(1-\hat{\mu}_{t}(x,0)\right)\Big{]}.$
 We denote the solution to the following optimization problem as $(\tilde{e}_{t+1}(\cdot),\dots,\tilde{e}_{T}(\cdot))$:
-$\min_{e_{t+1},\ldots,e_{T}}$ $\ \hat{\mathsf{P}}^{\text{forward}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)$ (7)
-subject to $\ \delta\leq e_{\ell}(\cdot)\leq 1-\delta,\quad\text{for }\ell=t+1,\ldots,T$
-$\ \hat{\mathsf{V}}^{\text{forward}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)+\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\big{(}\hat{\tau}_{t}(x)-\hat{\tau}_{t}\big{)}^{2}\leq C,$
+$\min_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}$ $\ \hat{\mathsf{P}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)$
+$\text{s.t. }\hat{\mathsf{V}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)+\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\big{(}\hat{\tau}_{t}(x)-\hat{\tau}_{t}\big{)}^{2}\leq C,$ (11)
 where $\hat{\tau}_{t}(x)=\hat{\mu}_{t}(x,1)-\hat{\mu}_{t}(x,0)$ and $\hat{\tau}_{t}=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\hat{\tau}_{t}(x)$.
-We then again set $\hat{e}^{*}_{t+1}(\cdot)=\tilde{e}_{t+1}(\cdot)$.
-To conclude our design strategy, as both optimization problems in Eqs  (6) and (7) are fully forward-looking, our proposed design strategies aim to proactively adjust the allocation strategies in future stages, ensuring that the data collected accounts for the future impacts due to outcome delays.
+We then again set $\hat{e}^{*}_{t+1}(1|\cdot)=1-\hat{e}^{*}_{t+1}(0|\cdot)=\tilde{e}_{t+1}(\cdot)$.
+As both optimization problems in (10) and (11) are fully forward-looking, our proposed design strategies aim to proactively adjust the allocation strategies in future stages, ensuring that the data collected accounts for the future impacts due to outcome delays.
 This approach contrasts with classical CARA designs that optimize trial objectives solely using historical information.
@@
 We then provide a variance estimator, thus enabling valid statistical inference:
-$\hat{\mathsf{V}}_{T}$ $=\sum_{x\in\mathcal{X}}\hat{p}_{T}(x)\Bigg{[}\frac{\hat{\sigma}_{T}^{2}(x,1)}{\frac{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=1,D_{i}+S_{i}\leq T)}}{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x)}}}+\frac{\hat{\sigma}_{T}^{2}(x,0)}{\frac{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=0,D_{i}+S_{i}\leq T)}}{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x)}}}+\big{(}\hat{\tau}_{T}(x)-\hat{\tau}\big{)}^{2}\Bigg{]}.$
-In the following sections, we will provide theoretical guarantees and simulation results to demonstrate the advantages of our proposed fully forward-looking CARA designs in the presence of delayed outcomes.
+$\hat{\mathsf{V}}_{T}$ $=\sum_{x\in\mathcal{X}}\hat{p}_{T}(x)\Bigg{[}\frac{\hat{\sigma}_{T}^{2}(x,1)}{\hat{\mathsf{e}}_{T}(x,1)}+\frac{\hat{\sigma}_{T}^{2}(x,0)}{\hat{\mathsf{e}}_{T}(x,0)}+\big{(}\hat{\tau}_{T}(x)-\hat{\tau}\big{)}^{2}\Bigg{]},\ \hat{\mathsf{e}}_{T}(x,a)=\frac{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq T)}}{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x)}}.$
+We summarize the analysis protocol below, where $z_{1-\frac{\alpha}{2}}$ denotes the $1-\frac{\alpha}{2}$ percentile of the standard normal distribution.
+Protocol for interim analysis and statistical inference
+Initial stage 1:
+   Enroll $n_{1}$ participants, and assign treatments with $\hat{e}_{1}^{*}(1|x)=\frac{1}{2}$.
+Interim analysis after stage $t$ has concluded:
+   Available data: $(S_{i},X_{i},A_{i},D_{i},Y_{i})$ for $i=1,2,\dots,N_{t}$;
+   Following (7) and (8), construct the delay distribution estimate $\hat{\rho}_{t}(d|x,a)$;
+   Following (9), construct the outcome distribution estimates $\hat{\sigma}_{t}(x,a)$ and $\hat{\mu}_{t}(x,a)$;
+   From either (10) or (11), solve the allocation for stage $t+1$: $\hat{e}_{t+1}^{*}(1|x)$.
+Treatment assignment for stage $t+1$:
+   Enroll $n_{t+1}$ participants, and assign treatments with $\hat{e}_{t+1}^{*}(1|x)$.
+Statistical inference after the final stage $T$ has concluded:
+   Available data: $(S_{i},X_{i},A_{i},D_{i},Y_{i})$ for $i=1,2,\dots,N$;
+   Construct the treatment effect estimator $\hat{\tau}_{T}$ and the variance estimator $\hat{\mathsf{V}}_{T}$;
+   Report the $(1-\alpha)\%$ confidence interval: $\hat{\tau}_{T}\pm z_{1-\frac{\alpha}{2}}\sqrt{\hat{\mathsf{V}}_{T}/N}$.
+In the following sections, we will provide theoretical guarantees and simulation evidence to demonstrate the advantages of our proposed fully forward-looking CARA designs in the presence of delayed outcomes.
 4 Theoretical investigation
 This section provides theoretical justifications for the proposed forward-looking CARA design.
-To begin, we present the semiparametric efficiency bound for estimating average treatment effects in the presence of response delay in multi-stage adaptive experiments.
+To begin, we present the semiparametric efficiency bound for treatment effect estimation in a multi-stage setting with delayed outcomes.
 We then develop a general result on the consistency and asymptotic normality of the estimated average treatment effect.
-Our theoretical investigation then proceeds by examining the specific setting considered in Section 3, where we extrapolate the unknown delay mechanism in each stage using a conservative rule.
-Encouragingly, we are able to justify Condition 1 in this setting under mild primitive assumptions.
-Finally, we observe that in an experiment with only finitely many stages, it is generally not possible to achieve the oracle-optimal allocation, as this would require perfect knowledge of the delay mechanism.
+Our theoretical investigation then proceeds by examining the specific setting considered in Section 3, where we extrapolate the estimated delay mechanism in each stage using a conservative rule.
+Finally, we observe that in an experiment with only finitely many stages, it is generally not possible to achieve the oracle allocation, as this would require perfect ex ante knowledge of the delay mechanism.
 Therefore, the last part of our theoretical investigation addresses the efficiency loss resulting from extrapolating the delay mechanism in our design method.
-Under conservative extrapolation, we show that the optimized asymptotic variance is guaranteed to weakly decrease as the experiment progresses.
-Our main results are developed under the general setting of multi-valued treatments, that is, there can be more than two treatment arms.
-Therefore, we will adopt a more general notation for the estimated treatment effects.
-As before, let $\hat{\mu}_{T}(x,a)$ denote the estimated mean potential outcome for treatment arm $a$ and subgroup $x$.
-Then the estimated subgroup treatment effect is $\hat{\tau}_{T}(x,a^{\prime},a)=\hat{\mu}_{T}(x,a^{\prime})-\hat{\mu}_{T}(x,a)$, and the estimated average treatment effect is $\hat{\tau}_{T}(a^{\prime},a)=\sum_{x\in\mathcal{X}}\hat{p}_{T}(x)\hat{\tau}_{T}(x,a^{\prime},a)$.
-4.1 Notation, assumptions and semiparametric efficiency bound
+4.1 Notation, assumptions and the efficiency bound
 We begin by recalling our adaptive experiment setting: the experiment consists of $T$ stages, and in each stage $t=1,2,\dots,T$, $n_{t}$ participants are enrolled.
-Let $N_{t}=\sum_{s=1}^{t}n_{s}$ represent the cumulative sample size at the end of stage $t$, with $N=N_{T}$ as the total sample size.
+Let $N_{t}=\sum_{s=1}^{t}n_{s}$ represent the cumulative sample size at the end of stage $t$, with $N=N_{T}$ being the total sample size.
 We use the notation $S_{i}=t$ to indicate that individual $i$ is enrolled in stage $t$.
-For each individual, we observe some baseline covariates, $X_{i}\in\mathcal{X}$.
-The observed outcome variable is denoted by $Y_{i}$, and it is related to the potential outcomes through the relation $Y_{i}=\sum_{a\in\mathcal{A}}\mathds{1}_{(A_{i}=a)}Y_{i}(a)$, where $A_{i}$ denotes the actual treatment assigned.
+For each individual, we observe some baseline covariates $X_{i}\in\mathcal{X}$.
+The observed outcome variable is denoted by $Y_{i}$, and it is related to the potential outcomes through $Y_{i}=A_{i}Y_{i}(1)+(1-A_{i})Y_{i}(0)$, where $A_{i}\in\mathcal{A}$ denotes the actual treatment assigned and $\mathcal{A}=\{0,1\}$.
+Moments of the potential outcomes are $\mu_{s}(x,a)=\mathbb{E}[Y_{i}(a)^{s}|X_{i}=x]$ and $\mu_{s}(a)=\mathbb{E}[Y_{i}(a)^{s}]$; we omit the subscript whenever $s=1$.
 Our first assumption imposes standard regularity conditions on the covariates and the potential outcomes.
 Assumption 1 (Covariates and potential outcomes).
-(i) The covariates and potential outcomes, $\{X_{i},\{Y_{i}(a):a\in\mathcal{A}\}\}$, are independently and identically distributed across $i=1,2,\dots,N$.
-(ii) The covariates have finite support, that is, $|\mathcal{X}|<\infty$.
-Let $p(x)=\mathbb{P}[X_{i}=x]$, then $\min_{x\in\mathcal{X}}p(x)>0$.
+(i) The covariates and potential outcomes, $(X_{i},Y_{i}(0),Y_{i}(1))$, are independent and identically distributed across $i=1,2,\dots,N$.
+(ii) The covariates have finite support (i.e., $|\mathcal{X}|<\infty$). Let $p(x)=\mathbb{P}[X_{i}=x]$, then $\min_{x\in\mathcal{X}}p(x)>0$.
 (iii) The potential outcomes have finite fourth moments: $\max_{x\in\mathcal{X}}\max_{a\in\mathcal{A}}\mu_{4}(x,a)<\infty$.
-The next assumption addresses the delay mechanism.
-Recall that the delay is represented by the random variable $D_{i}$, which indicates the number of stages after which the experimenter can observe $Y_{i}$.
+The next assumption addresses the delay mechanism, where we adopt the notation $D_{i}$ to indicate the number of stages after which the experimenter can observe $Y_{i}$.
 By our definition, $Y_{i}$ is observed at the end of stage $t$ if and only if $D_{i}+S_{i}\leq t$.
+Also recall that the delay distribution is represented by $\rho(d|x,a)=\mathbb{P}[D_{i}\leq d|X_{i}=x,A_{i}=a]$.
 Assumption 2 (Delay mechanism).
 […]
-Finally, as we consider multi-stage experiment settings, we are in an asymptotic regime in which $T$ is fixed, and $n_{t}\to\infty$.
+(ii) $\min_{x\in\mathcal{X}}\min_{a\in\mathcal{A}}\rho(0|x,a)>0$.
+Finally, we consider multi-stage experiment settings.
 Assumption 3 (Asymptotic regime).
 The total number of stages, $T$, is fixed, and for all $t$, $n_{t}/N\to r_{t}>0$.
-We are now ready to present the semiparametric efficiency bound for average treatment effect estimation in the oracle setting, where there is perfect knowledge of the delay mechanism, and treatments are randomly assigned conditional on the covariates.
-This efficiency bound serves as the foundation and starting point for our proposed adaptive experimental design, as our algorithm minimizes a feasible version of the efficiency bound, constructed using accrued information.
+We are now ready to present the semiparametric efficiency bound for average treatment effect estimation in the oracle setting, where one has perfect knowledge of the delay mechanism and treatments are randomly assigned conditional on the covariates.
+This efficiency bound serves as the foundation and starting point for our proposed adaptive experimental design, as our algorithm minimizes a feasible version of the efficiency bound constructed using accrued information.
 Theorem 1 (Semiparametric efficiency bound).
 Let Assumptions 13 hold.
-In addition, assume (i) $S_{i}$ are independently and identically distributed with $\mathbb{P}[S_{i}=t]=r_{t}$; (ii) the treatment assignment probabilities, $e_{t}(a|x)=\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t]$, are bounded away from 0 and 1.
-Then the efficient influence function for estimating $\mu_{1}(a)$ is
-$\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a)$ $=\frac{\mathds{1}_{(A_{i}=a,D_{i}+S_{i}\leq T)}}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|X_{i})\rho(T-t|X_{i},a)}\Big{(}Y_{i}(a)-\mu_{1}(X_{i},a)\Big{)}+\mu_{1}(X_{i},a)-\mu_{1}(a).$
-In addition, the efficiency influence function for estimating $\tau(a^{\prime},a)$ is
-$\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a^{\prime},a)\ =\ \psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a^{\prime})-\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a)$
-$=$ $\frac{\mathds{1}_{(A_{i}=a^{\prime},D_{i}+S_{i}\leq T)}}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a^{\prime}|X_{i})\rho(T-t|X_{i},a^{\prime})}\Big{(}Y_{i}(a^{\prime})-\mu_{1}(X_{i},a^{\prime})\Big{)}$
-$\qquad\qquad-\frac{\mathds{1}_{(A_{i}=a,D_{i}+S_{i}\leq T)}}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|X_{i})\rho(T-t|X_{i},a)}\Big{(}Y_{i}(a)-\mu_{1}(X_{i},a)\Big{)}+\tau(X_{i},a^{\prime},a)-\tau(a^{\prime},a),$
+In addition, assume (i) $S_{i}$ are independent and identically distributed with $\mathbb{P}[S_{i}=t]=r_{t}$; (ii) the treatment assignment probabilities, $e_{t}(a|x)=\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t]$, are bounded away from 0 and 1.
+Then the efficient influence function for estimating $\mu(a)$ is
+$\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a)$ $=\frac{\mathds{1}_{(A_{i}=a,D_{i}+S_{i}\leq T)}}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|X_{i},a)e_{t}(a|X_{i})}\Big{(}Y_{i}(a)-\mu(X_{i},a)\Big{)}+\mu(X_{i},a)-\mu(a),$
+where $a\in\mathcal{A}=\{0,1\}$. In addition, the efficient influence function for estimating $\tau$ is
+$\psi(X_{i},A_{i},D_{i},Y_{i},S_{i})\ =\ \psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|1)-\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|0),$
 leading to the semiparametric efficiency bound:
-$\mathsf{V}=\sum_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,a^{\prime})}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a^{\prime}|x)\rho(T-t|x,a^{\prime})}$ $+\frac{\sigma^{2}(x,a)}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}+\big{(}\tau(x,a^{\prime},a)-\tau(a^{\prime},a)\big{)}^{2}\Bigg{]}.$
+$\mathsf{V}=\sum_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,1)}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|x,1)e_{t}(1|x)}$ $+\frac{\sigma^{2}(x,0)}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|x,0)e_{t}(0|x)}+\big{(}\tau(x)-\tau\big{)}^{2}\Bigg{]}.$
 We note that while the previous theorem is developed in the context of discrete covariates (with the efficiency bound expressed as a summation over the probability mass function), it can be generalized to accommodate, for example, continuously distributed covariates (c.f., equation 1).
 4.2 Asymptotic properties of treatment effects estimates
-In this subsection, we demonstrate that the treatment effect estimator is consistent and admits an asymptotically normal distribution.
-The main conclusions of this section rely on a high-level condition regarding the consistency of the optimized treatment allocation rule, making these conclusions applicable to a wide range of CARA and a broad class of frequentist adaptive experimental design settings (Hu and Zhang, 2004; Baldi Antognini and Zagoraiou, 2015; Hu et al., 2015).
-In line with existing literature, the consistency of optimized treatment allocation rules is often needed to establish statistical guarantees of various adaptive designs.
-We recall that the treatment assignment probabilities are denoted by $\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t,\mathcal{H}_{t-1}]=\hat{e}^{*}_{t}(a|x)\in\mathcal{H}_{t-1}$, which are allowed to depend on historical data.
-We begin with a high-level condition, which will be verified in the next subsection for our proposed forward-looking CARA design.
+In this subsection, we demonstrate that the treatment effect estimator is consistent and admits an asymptotic normal distribution.
+The main conclusions of this section rely on a high-level condition regarding the convergence of the optimized treatment allocation rule, making these conclusions applicable to a wide range of CARA and a broad class of frequentist adaptive experimental design settings (Hu and Zhang, 2004; Antognini and Zagoraiou, 2015; Hu, Zhu and Hu, 2015).
+We begin with this high-level condition, which will be verified in the next subsection for our proposed forward-looking CARA design.
 Condition 1 (Convergence of the optimized treatment allocation).
 There exists nonrandom $e_{t}(\cdot|\cdot)$, such that $\hat{e}^{*}_{t}(a|x)=e_{t}(a|x)+o_{\mathrm{p}}(1)$ for all $a\in\mathcal{A}$ and all $x\in\mathcal{X}$.
-We then present a useful lemma, which characterizes an asymptotically linear representation of the estimated mean potential outcomes.
-Lemma 1 (Asymptotically linear representation for estimated mean potential outcomes).
+The next theorem characterizes the asymptotic normality of both the estimated subgroup treatment effects and the estimated average treatment effect.
+Theorem 2 (Asymptotic normality of estimated treatment effects).
 Let Assumptions 13 hold. Then under Condition 1,
-$\sqrt{N}\big{(}\hat{\mu}_{1,T}(x,a)-\mu_{1}(x,a)\big{)}$
-$\quad=\frac{1}{p(x)\sum_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq T)}\big{(}Y_{i}-\mu_{1}(x,a)\big{)}+o_{\mathrm{p}}(1).$
-The next theorem characterizes the asymptotic normality of both the estimated subgroup treatment effects and the estimated average treatment effect.
-Theorem 2 (Asymptotic normality of estimated subgroup treatment effects).
-Let Assumptions 13 hold. Then under Condition 1,
-$\sqrt{N}\big{(}\hat{\tau}_{T}(x,a^{\prime},a)-\tau(x,a^{\prime},a)\big{)}\overset{\mathrm{d}}{\to}\mathcal{N}\Big{(}0,\frac{1}{p(x)}\Big{(}\frac{\sigma^{2}(x,a^{\prime})}{\sum_{t=1}^{T}r_{t}e_{t}(a^{\prime}|x)\rho(T-t|x,a^{\prime})}+\frac{\sigma^{2}(x,a)}{\sum_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}\Big{)}\Big{)},$
-and
-$\sqrt{N}\big{(}\hat{\tau}_{T}(a^{\prime},a)-\tau(a^{\prime},a)\big{)}\overset{\mathrm{d}}{\to}\mathcal{N}(0,\mathsf{V}),$
-where $\mathsf{V}$ is defined in Theorem 1.
-Before discussing statistical inference, we compare this theorem to the semiparametric efficiency bound established in the previous subsection.
-Specifically, while the two results, Theorems 1 and 2, yield the same variance formula, they are conceptually very different.
-The semiparametric efficiency bound is derived under a fixed treatment assignment regime, which we then use to guide our experimental design algorithm.
-Specially, our algorithm aims to minimize this efficiency bound.
-However, from an ex ante perspective, it may be unclear why this is a useful exercise, as it is not immediately apparent how to achieve this efficiency bound.
-Encouragingly, Theorem 2 helps “close the loop” by showing that, under mild regularity conditions, the average treatment effect estimator achieves the efficiency bound, thereby justifying our design objective.
-To conclude this subsection, the result below establishes valid statistical inference, where the variance estimator, $\hat{\mathsf{V}}_{T}$, is defined at the end of Section 3.
+$\sqrt{N}\big{(}\hat{\tau}_{T}(x)-\tau(x)\big{)}\overset{\mathrm{d}}{\to}\mathcal{N}\Big{(}0,\frac{1}{p(x)}\Big{(}\frac{\sigma^{2}(x,1)}{\sum_{t=1}^{T}r_{t}\rho(T-t|x,1)e_{t}(1|x)}+\frac{\sigma^{2}(x,0)}{\sum_{t=1}^{T}r_{t}\rho(T-t|x,0)e_{t}(0|x)}\Big{)}\Big{)},$
+and $\sqrt{N}\big{(}\hat{\tau}_{T}-\tau\big{)}\overset{\mathrm{d}}{\to}\mathcal{N}(0,\mathsf{V})$, where $\mathsf{V}$ is defined in Theorem 1.
+It is helpful to compare this theorem to the semiparametric efficiency bound established in the previous subsection.
+While the two results, Theorems 1 and 2, yield the same variance formula, they are conceptually very different.
+The semiparametric efficiency bound is derived under a fixed treatment assignment regime, which we then use to guide our design algorithms—either by minimizing the efficiency bound for power improvement or by maximizing failure reduction subject to a variance upper bound.
+However, from an ex ante perspective, it may not be immediately clear why this is a useful exercise, as it is not obvious how the bound could actually be achieved.
+Encouragingly, Theorem 2 helps close the loop by showing that the asymptotic variance of the estimated treatment effect matches the bound in Theorems 1 under mild regularity conditions, thereby justifying our design objective.
+That said, it is worth clarifying that, in general, one cannot achieve the optimized oracle efficiency bound, an issue we discuss further below.
+To conclude this subsection, the following result establishes valid statistical inference, where the variance estimator $\hat{\mathsf{V}}_{T}$ is defined at the end of Section 3.
 Theorem 3 (Statistical inference).
@@
 These results build on the high-level condition that the empirical treatment allocation, $\hat{e}_{t}^{*}(\cdot)$, will converge in large samples (Condition 1).
-In this subsection, we further investigate the theoretical properties of the proposed experimental design—specifically, solved from (6)—and show that it satisfies Condition 1, thereby tying all loose ends.
+In this subsection, we further investigate the theoretical properties of the proposed experimental design, specifically, solved from (10), and show that it satisfies Condition 1, thereby tying all loose ends.
 To give a road map, we will first show that the asymptotic analogs of our design algorithm have unique solutions.
-In this process, we will carefully distinguish the oracle problem of minimizing the semiparametric efficiency bound (which requires perfect knowledge of the delay mechanism) from the problem of sequential optimization using only accrued knowledge (which requires extrapolating the delay mechanism).
+In this process, we will carefully distinguish the oracle problem of minimizing the semiparametric efficiency bound (which requires perfect knowledge of the delay mechanism) from the problem of sequential optimization using only accrued knowledge (which requires extrapolating the estimated delay mechanism).
 Next, we show that the empirically optimized treatment allocation converges to some large-sample limit.
@@
 This allows us to provide simple and easy-to-interpret conditions, hence avoiding lengthy discussions.
-Despite our focus on one particular approach in this subsection, we show in our simulations that all three methods (conservative, optimistic, and neutral) deliver valid inference for the desired causal effects and more effectively achieves the design objectives compared to classical approaches.
-To start, we recall that $\mathsf{V}=\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ denotes the efficiency bound. Its minimizers will be denoted by
-$\{e_{t}^{*}(\cdot),1\leq t\leq T\}=\operatorname*{argmin}_{e_{1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot)),\ \text{subject to }e_{t}(a|\cdot),e_{t}(a^{\prime}|\cdot)\geq\delta.$
-As the semiparametric efficiency bound is specific to two treatment values, $a$ and $a^{\prime}$, we impose the constraints that the treatment probabilities in these two arms are bounded from below.
-However, it is worth mentioning that one can easily incorporate additional constraints.
-As we have discussed, the oracle optimal solution $\{e_{t}^{*}(\cdot),1\leq t\leq T\}$ is not achievable in general, as it builds on perfect knowledge of the delay mechanism.
+Another appealing property of the conservative approach is that it can be understood from a “minimax” perspective, where we sequentially minimize the asymptotic variance under the worst-case scenario of the delay mechanism.
+Despite our focus on one particular approach in this subsection, we show in our simulations that all three methods (conservative, optimistic, and neutral) deliver valid inference.
+To start, we recall that $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ is the objective function in (3).
+Its minimizers will be denoted by $e_{t}^{*}(a|x)$ for $a\in\{0,1\}$, $x\in\mathcal{X}$, and $t=1,2,\dots,T$.
+As we have discussed, the oracle optimal solution is not achievable in general, as it builds on perfect knowledge of the delay mechanism.
 In any realistic experimental setting, however, this information can only be learned sequentially.
-In fact, even we set $n_{t}=\infty$, part of the delay mechanism can still remain unknown for any $t$: to be even more precise, one typically only learn $\rho(d|\cdot)$ for $d<t$ using information at the end of stage $t$.
+In fact, even we set $n_{t}=\infty$, part of the delay mechanism can still remain unknown for any $t$: to be even more precise, one typically only learn $\rho(d|x,a)$ for $d<t$ using information at the end of stage $t$.
 For this reason, we have introduced feasible and concrete ways to extrapolate the delay mechanism.
 Following our discussion in Section 2, we define
-$\rho^{\dagger}_{t}(d|\cdot)$ $=\min\big{\{}\rho(d|\cdot),\rho(t-1|\cdot)\big{\}}\ \text{for $d=1,2,\dots$}.$
-Alternatively, we write $\rho^{\dagger}_{t}(d|\cdot)=\rho(d|\cdot)$ for $d<t$, and $\rho^{\dagger}_{t}(d|\cdot)=\rho(t-1|\cdot)$ for $d=t,t+1,\dots$.
+$\rho^{\dagger}_{t}(d|x,a)$ $=\min\big{\{}\rho(d|x,a),\rho(t-1|x,a)\big{\}}.$
 It is helpful to compare this definition with Assumption 2.
-There, we assumed that the delay mechanism is time-homogeneous, so that $\rho(\cdot)$ is not indexed by $t$.
-On the other hand, $\rho^{\dagger}_{t}(\cdot)$ arises due to our imperfect knowledge about the delay mechanism and the conservative extrapolation employed, and as a result it will depend on a specific stage.
-It is not the true delay mechanism, but it is the asymptotic analogue of $\hat{\rho}_{t}(\cdot)$, which we used in our optimization algorithm.
-To study convergence property of $\hat{e}^{*}_{t}(\cdot)$, we will make a definition first.
-Let $e^{\dagger}_{1}(a|\cdot)=e^{\dagger}_{1}(a^{\prime}|\cdot)=1/2$.
-Then for $t=2,3,\dots,T$, define recursively that
-$e^{\dagger}_{t}(\cdot)=\operatorname*{argmin}_{e_{t}(\cdot)}\min_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t}^{\dagger}(e_{t}(\cdot),e_{t+1}(\cdot),\dots,e_{T}(\cdot)),\ \text{subject to }e_{t}(a|\cdot),e_{t}(a^{\prime}|\cdot)\geq\delta,$
+There, we assumed that the delay mechanism is time-homogeneous, so that $\rho(d|x,a)$ is not indexed by $t$.
+On the other hand, $\rho^{\dagger}_{t}(d|x,a)$ arises due to our imperfect knowledge about the delay mechanism and the conservative extrapolation employed, and as a result it will depend on a specific stage.
+It is not the true delay mechanism, but it is the asymptotic analogue of $\hat{\rho}_{t}(d|x,a)$, which we used in our optimization algorithm.
+To study convergence property of $\hat{e}^{*}_{t}(a|x)$, we will make a definition first.
+Let $e^{\dagger}_{1}(0|x)=e^{\dagger}_{1}(1|x)=1/2$.
+Then for $t=1,2,\dots,T$, define recursively that
+$e^{\dagger}_{t+1}(1|\cdot)=1-e^{\dagger}_{t+1}(0|\cdot)=\operatorname*{argmin}_{e_{t+1}(\cdot)}\min_{e_{t+2}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot)),$
 where
-$\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,a^{\prime})}{\sum\limits_{\ell=1}^{t}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a^{\prime})e_{\ell}^{\dagger}(a^{\prime}|x)+\sum\limits_{\ell=t+1}^{T}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a^{\prime})e_{\ell}(a^{\prime}|x)}$
-$+\frac{\sigma^{2}(x,a)}{\sum\limits_{\ell=1}^{t}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a)e_{\ell}^{\dagger}(a|x)+\sum\limits_{\ell=t+1}^{T}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a)e_{\ell}(a|x)}\Bigg{]}.$
-Our first main result is to show that both $e^{*}_{t}(\cdot)$ and $e^{\dagger}_{t}(\cdot)$ are unique. To this end, the following assumption.
+$\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,1)}{\sum\limits_{\ell=1}^{t}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,1)e_{\ell}^{\dagger}(1|x)+\sum\limits_{\ell=t+1}^{T}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,1)e_{\ell}(x)}$
+$+\frac{\sigma^{2}(x,0)}{\sum\limits_{\ell=1}^{t}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,0)e_{\ell}^{\dagger}(0|x)+\sum\limits_{\ell=t+1}^{T}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,0)(1-e_{\ell}(x))}\Bigg{]}.$
+Our first main result is to show that both $e_{t}^{*}(a|x)$ and $e_{t}^{\dagger}(a|x)$ are unique. To this end, we make the following assumption.
 Assumption 4 (Variation in delay mechanism).
-For any $d\neq d^{\prime}$, ${\rho(d|x,a^{\prime})}/{\rho(d|x,a)}\neq{\rho(d^{\prime}|x,a^{\prime})}/{\rho(d^{\prime}|x,a)}$.
-We remark that this assumptions helps establish the uniqueness of $e^{*}_{t}(\cdot)$.
-However, it is still possible to have ${\rho_{t}^{\dagger}(d|x,a^{\prime})}/{\rho_{t}^{\dagger}(d|x,a)}={\rho_{t}^{\dagger}(d^{\prime}|x,a^{\prime})}/{\rho_{t}^{\dagger}(d^{\prime}|x,a)}$.
+For any $d\neq d^{\prime}$, ${\rho(d|x,1)}/{\rho(d|x,0)}\neq{\rho(d^{\prime}|x,1)}/{\rho(d^{\prime}|x,0)}$.
+We remark that this assumptions helps establish the uniqueness of $e_{t}^{*}(a|x)$.
+However, it is still possible to have ${\rho_{t}^{\dagger}(d|x,1)}/{\rho_{t}^{\dagger}(d|x,0)}={\rho_{t}^{\dagger}(d^{\prime}|x,1)}/{\rho_{t}^{\dagger}(d^{\prime}|x,0)}$.
 Fortunately, this is easily resolved by a tie-breaking rule.
@@
 Let Assumptions 14 hold.
-Then both $e_{t}^{*}(\cdot)$ and $e_{t}^{\dagger}(\cdot)$ are unique.
-In addition, Condition 1 holds with $\hat{e}_{t}^{*}(\cdot)\overset{\mathrm{p}}{\to}e_{t}^{\dagger}(\cdot)$.
+Then for $a\in\{0,1\}$, $x\in\mathcal{X}$, and $t=1,2,\dots,T$, both $e_{t}^{*}(a|x)$ and $e_{t}^{\dagger}(a|x)$ are unique.
+In addition, Condition 1 holds with $\hat{e}_{t}^{*}(a|x)\overset{\mathrm{p}}{\to}e_{t}^{\dagger}(a|x)$.
 Before closing this section, we provide two insights regarding the variance minimization problem.
-The first result suggests that as more information about the delay mechanism becomes available in the adaptive experiment, the optimized asymptotic variance never increases.
-While this result may seem quite natural, we remark that it does rely on the specific extrapolation we employed for imputing the delay mechanism, and to be specific, the conservative method we discussed earlier.
+The first result suggests that as more information about the delay mechanism becomes available in the adaptive experiment, the optimized asymptotic variance always (weakly) decreases.
+While this result may seem natural, we note that it relies on the use of conservative extrapolation.
+As discussed earlier, conservative extrapolation can be understood from a minimax perspective, where the asymptotic variance is sequentially minimized under the worst-case scenario of the delay mechanism.
 The second result provides a bound on efficiency loss due to imperfect knowledge about the delay mechanism.
@@
 Theorem 5 (Optimized variance and bound on efficiency loss).
-Let Assumptions 14 hold. Define $\mathsf{V}^{*}$ as the minimized oracle asymptotic variance. Then for all $t=1,2,\dots,T-1$,
-$\min_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))\ \geq$ $\min_{e_{t+2}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t+1}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))\ \geq\ \mathsf{V}^{*}.$
-Denote by $\mathsf{V}$ the asymptotic variance of the estimated average treatment effect. Then $\mathsf{V}\leq\mathsf{V}^{*}+O(1/T)$.
+Let Assumptions 14 hold.
+Define $\mathsf{V}^{*}$ as the minimized oracle asymptotic variance, and $\mathsf{V}$ to be the asymptotic variance of the estimated average treatment effect.
+Then (i) for all $1\leq s<t\leq T-1$,
+$\mathsf{V}^{*}\ \leq\ \mathsf{V}\ \leq\ \min_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))\ \leq\ \min_{e_{s+1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{s}^{\dagger}(e_{s+1}(\cdot),\dots,e_{T}(\cdot));$
+and (ii) for any $\epsilon>0$ and $d_{\epsilon}$ with $\min_{x\in\mathcal{X},a\in\mathcal{A}}\rho(d_{\epsilon}|x,a)\geq 1-\epsilon$, $\mathsf{V}\leq\mathsf{V}^{*}+C(d_{\epsilon}T^{-1}+\epsilon)$, where $C$ does not depend on $T$, $\epsilon$ or $d_{\epsilon}$.
 5 Synthetic case study
-To evaluate the performance of the proposed forward-looking CARA design, we conduct a synthetic case study based on the randomized experiment reported in Fahey et al.
+To evaluate the performance of the proposed forward-looking CARA design, we conduct simulation studies using a data generation mechanism calibrated from the study of Fahey, Njau, Katabaro, Mfaume, Ulenga, Mwenda et al.
 (2020).
-Through this synthetic case study, we first demonstrate that both of the two forward-looking CARA designs (Section 3) deliver valid statistical inference (Table 1), supporting the theoretical result established in Theorem 2.
-Second, numerical evidence suggests that our fully forward-looking CARA design with the power maximization objective achieves a higher estimation efficiency compared to both the Neyman allocation and the complete randomization design in the presence of delayed responses (Figures 5 and 6), further validating the theoretical results in Theorem 5.
-Finally, compared to both the ethical design (Rosenberger et al., 2001) and complete randomization, our fully forward-looking CARA design can provide greater failure reduction, demonstrating its effectiveness in improving participant welfare while maintaining statistical power (Figure 5).
-The original experiment conducted by Fahey et al.
-(2020) aims to evaluate the effectiveness of cash transfers on retention in care among people living with HIV (PLHIV), using viral load as the primary response variable.
-The experiment participants were adults aged 18 years or older diagnosed with HIV, who were randomized to receive cash transfers (treatment arm) or no cash transfers (control arm).
+Their original experiment aims to evaluate the effectiveness of cash transfers on retention in care among people living with HIV (PLHIV), using viral load as the primary response variable.
+The experiment participants were adults aged 18 years or older diagnosed with HIV, who were randomized to receive cash transfers (treatment arm, $A_{i}=1$) or no cash transfer (control arm, $A_{i}=0$).
+The original study involved two treatment arms with different levels of cash incentives, which are combined into one group in our synthetic case study.
 The response variable, viral load, was assessed six months after enrollment.
 Although the original study was not conducted in an adaptive fashion, adaptation can be applied since the participants were enrolled sequentially.
-However, a challenge is that the primary responses of the participants were collected with delays, with some responses collected longer than six months after the treatment assignment.
-As illustrated in Example 1, the empirically observed delay mechanism exhibits significant heterogeneity across arms and covariate strata.
-Our fully forward-looking CARA design can be utilized in such scenarios to help the experiment more effectively achieve its design goals.
-In this synthetic case study, we conduct a simulation experiment in which the underlying data-generating process — including the response delay mechanism, covariate distribution, and the (conditional) potential outcomes distributions — is calibrated with estimates from the (Fahey et al., 2020) study.
-We set $n_{t}=100$ for $t=1,\ldots,T$, and $T$ ranges in $\{4,6,8\}$.
-The primary response variable is the viral load measured at the end of each stage.
-We consider two different transformations of the original viral load measurements, which shall be illustrated in the two simulation setups below.
-Delays in collecting the response variable are observed, and as discussed in Example 1, the delay mechanism depends on the arm and the biological sex of the participant.
-We consider two treatment arms: the control arm $A_{i}=0$, defined as “received no cash transfer”, and the treatment arm, $A_{i}=1$, defined as “received cash transfer.” Biological sex is included as the covariate $X_{i}$, where $X_{i}=1$ and 00 represent “male” and “female,” respectively.
-From the original experimental data, the distribution of biological sex consists of 36% male and 64% female, which we adopt as the data-generating process of $X_{i}$ in our Monte Carlo exercise.
-Specifically, we generate the covariate $X_{i}$ and treatment $A_{i}$ as follows:
-$X_{i}$ $\sim\text{Bernoulli}(0.36),\ i=1,\ldots,N,$
-$A_{i}|X_{i}=x,S_{i}=t$ $\sim\text{Bernoulli}(e_{t}(1|x)),\ t=1,\ldots,T,$
-where $e_{t}(1|x)$ entails different designs we aim to compare.
-In what follows, we shall illustrate the design strategies in comparison under two sets of simulation setups tailored to each design objective:
+However, a challenge is that the primary responses of the participants were collected with delays, with some responses collected longer than six months after the treatment assignment; see Figure 1.
+Our synthetic case study conducts a simulation experiment in which the underlying data-generating process (delay mechanism, covariate distribution, and potential outcome distributions) is calibrated with estimates from the original study.
+For the first set of simulation results, we consider biological sex as the only covariate $X_{i}$, leading to two subgroups: $X_{i}=\mathtt{M}$ and $X_{i}=\mathtt{F}$ for male and female participants.
+From the original data, the distribution of biological sex consists of 36% male and 64% female, which we adopt as the data-generating process of $X_{i}$ in our Monte Carlo exercise.
+We set $n_{t}=100$ for $t=1,\ldots,T$ with $T\in\{4,6,8\}$.
+We report simulation evidence separately for the two design objectives in (3) and (4), which require different transformation of the outcome variable.
 Setup 1: Power maximization.
-For the power maximization design objective in Eq (2), we compare four designs: (1) the fully forward-looking CARA design for power maximization (Section 3); (2) the fully forward-looking CARA design further enhanced by the doubly adaptive biased coin (DBCD) method (Hu and Zhang, 2004; Tymofyeyev et al., 2007); (3) complete randomization; (4) treatment allocation based on Neyman allocation.
-Furthermore, in HIV research, it is standard practice to apply a log transformation to raw viral load measurements, given their wide range.
-Consequently, the response variable in this simulation setup is defined as: $\log(\texttt{viral load})$.
-We generate the response variable as
-$Y_{i}|A_{i}=a,X_{i}=x\sim\mathcal{N}\left(\mu(x,a),\ \sigma^{2}(x,a)\right)$
-to mimic the distribution of $\log(\texttt{viral load})$ in the original data, where for women $\mu(0,1)=2.50$, $\mu(0,0)=2.98$, $\sigma(0,1)=0.36$, $\sigma(0,0)=2.06$, and for men $\mu(1,1)=2.47$, $\mu(1,0)=2.72$, $\sigma(1,1)=0.82$, and $\sigma(1,0)=0.31$.
-Lastly, to evaluate the power of different designs under the power maximization objective, we vary the magnitude of ATE in our data-generating process by perturbing $\mu(1,0)$, such that the ATE $\tau(1,0)=0.33+\delta$, where $\delta\in\{-0,33,\ldots,0,\ldots,0.33\}$.
+For the power maximization design objective, we compare four design methods: (i) our proposed design; (ii) our design enhanced by the doubly adaptive biased coin design (DBCD) (Hu and Zhang, 2004; Tymofyeyev, Rosenberger and Hu, 2007); (iii) complete randomization; (iv) Neyman allocation.
+We note that DBCD is a response-adaptive randomization design that more robustly adjusts treatment allocation toward the optimal allocation dynamically.
+Since our approach involves assigning treatments across multiple covariate strata, we apply the DBCD design separately within each stratum.
+We also follow the convention in the literature and use the natural logarithm of the viral load as the outcome variable.
+Following the trial data, the conditional mean and standard deviation of the potential outcomes are $\mu(\mathtt{F},1)=2.50$, $\mu(\mathtt{F},0)=2.98$, $\sigma(\mathtt{F},1)=0.36$, $\sigma(\mathtt{F},0)=2.06$ for the female subgroup, and $\mu(\mathtt{M},1)=2.47$, $\mu(\mathtt{M},0)=2.72$, $\sigma(\mathtt{M},1)=0.82$, $\sigma(\mathtt{M},0)=0.31$ for the male subgroup.
+Given the calibrated model primitives, the average treatment effect is $\tau=-0.33$.
+To evaluate the power of different designs under the power maximization objective, we also vary the magnitude of the ATE in our data-generating process by perturbing $\mu(\mathtt{M},0)$, such that the perturbed ATE takes values between 0 and $0.66$.
 Setup 2: Failure reduction.
-For the failure reduction design objective in Eq (3), we compare four designs: (1) the proposed fully forward-looking CARA design for failure reduction; (2) the proposed design enhanced by the DBCD method (Hu and Zhang, 2004); (3) the ethical design introduced by Rosenberger et al.
-(2001); (4) complete randomization.
-In this simulation setup, since the goal is to reduce the percentage of failures, we transform the original viral load measurements into a binary response variable, indicating whether the participant has achieved viral suppression.
-Specifically, if a participant’s viral load is less than $1,000$ copies per mL, we set $Y_{i}=1$, representing viral suppression.
-Otherwise, $Y_{i}$ is set to 00 (no viral suppression).
-We generate the binary response variable as
-$Y_{i}|A_{i}=a,X_{i}=x\sim\text{Bernoulli}(\mu(x,a)),\ i=1,\ldots,N,$
-where $\mu(0,1)=0.78$ and $\mu(0,0)=0.57$ represent the expected probabilities of viral suppression for females in the treatment and control arms, respectively.
-Similarly, $\mu(1,1)=0.84$ and $\mu(1,0)=0.63$ represent the expected probabilities of viral suppression probabilities for males in the treatment and control arms, respectively.
-Lastly, to evaluate the power of different designs under the failure reduction objective, we vary the magnitude of ATE in our data-generating process.
-For Setup 2, we vary the ATE value by perturbing $\mu(1,0)$, such that the ATE $\tau(1,0)=0.21+\delta$, where $\delta\in\{-0,21,\ldots,0,\ldots,0.21\}$.
-In our synthetic case study, we incorporate both arm- and covariate-specific outcome delays following empirical estimates obtained in the original experimental data.
-We generate the delay variable $D_{i}$ from multinomial distribution:
-$D_{i}|A_{i}=1,X_{i}=1$ $\sim\text{Multinomial}\left(1,\bm{\rho}(\cdot|1,1)\right),\ D_{i}|A_{i}=1,X_{i}=0\sim\text{Multinomial}\left(1,\bm{\rho}(\cdot|1,0)\right),$
-$D_{i}|A_{i}=0,X_{i}=1$ $\sim\text{Multinomial}\left(1,\bm{\rho}(\cdot|0,1)\right),\ D_{i}|A_{i}=0,X_{i}=0\sim\text{Multinomial}(1,\bm{\rho}(\cdot|0,0)),$
-where the first row represents the response delay mechanism for males and females under the treatment arm, while the second row represents the delay mechanism for males and females under the control arm.
-We slightly abuse the notation to define $\bm{\rho}(\cdot|a,x):=(\mathbb{P}(D_{i}=0|a,x),\ldots,\mathbb{P}(D_{i}=T-1|a,x))^{\intercal}$, where each element represents the probability that the response of subject $i$ is delayed by exactly $d$ stages, for $d=0,\ldots,T-1$.
-When $T=4$, the parameter values for the delay mechanisms are as follows: $\bm{\rho}(\cdot|1,1)=(0.64,0.18,0.07,0.03)^{\intercal}$, $\bm{\rho}(\cdot|1,0)=(0.63,0.18,0.05,0.02)^{\intercal}$, $\bm{\rho}(\cdot|0,1)=(0.55,0.23,0.10,0.02)^{\intercal}$, and $\bm{\rho}(\cdot|0,0)=(0.54,0.11,0.21,0.01)^{\intercal}$.
-We defer the details of the delay mechanism generation for $T=6$ and $T=8$ to the Supplementary Materials.
-To generate delayed responses, at each Stage $t$, we first generate responses for all $n_{t}$ subjects.
-Based on the delay mechanism, we then randomly select a subset of subjects with $(A_{i}=a,X_{i}=x)$ and label their responses as NA, indicating that their responses will be delayed to be observed.
-For both simulation setups, we generate 500 Monte Carlo samples.
-To initialize the experiment, we set the treatment allocation to be $1/2$ in each covariate stratum in Stage 1 for all the designs.
-For Setup 1, we shall compare the variance of the ATE estimator under different design strategies in Figure 5 (A) – (C).
-For Setup 2, we shall compare the failure reduction rate under different design strategies in Figure 5 (D) – (E).
-For Setup 1 and Setup 2, we consider the conservative experimenter’s expectations about future delays.
-We defer the simulation results for neutral and optimistic experimenter’s expectations to Supplementary Materials.
-For both setups, we also present the coverage probabilities (Table 1) and the power curve (Figure 6) to evaluate the statistical validity.
-Figure 5: The variance of the ATE estimator under the power maximization design objective (A) – (C) and the failure reduction rate under the failure reduction design objective (D) – (F) across different design strategies.
-Figure 6: The power curves under the (A) power maximization design objective and the (B) failure reduction design objective across various design strategies.
+For the failure reduction objective, we again compare across four designs: (i) the proposed method; (ii) our proposed design again enhanced by DBCD; (iii) the ethical design proposed by Rosenberger, Stallard, Ivanova, Harper and Ricks (2001); (iv) complete randomization.
+In this simulation setup, we transform the original viral load measurements into a binary variable, indicating whether the participant has achieved viral suppression; specifically, if a participant’s viral load is below $1,000$ copies per mL, we set $Y_{i}=1$ (viral suppression).
+Conditional means of the potential outcomes are $\mu(\mathtt{F},1)=0.78$, $\mu(\mathtt{F},0)=0.57$, $\mu(\mathtt{M},1)=0.84$ and $\mu(\mathtt{M},0)=0.63$.
+The average treatment effect is $\tau=0.21$.
+To evaluate the power of different designs under the failure reduction objective, we vary the magnitude of ATE in our data-generating process by perturbing $\mu(\mathtt{M},0)$, such that the perturbed ATE takes values between 0 and $0.42$.
+As we discussed in Section 1 and illustrated in Figure 1, the outcomes in the original data exhibit both arm- and covariate-dependent delays, which we also incorporate in our simulation setup.
+To be precise, the delay variable is generated conditionally according a multinomial distribution: $D_{i}|(X_{i}=x,A_{i}=a)\sim\text{Multinomial}\left(1,\newline \bm{\rho}(x,a)\right)$.
+Here, we slightly abuse the notation to define $\bm{\rho}(x,a)=(\mathbb{P}(D_{i}=0|x,a),\ldots,\mathbb{P}(D_{i}=T-1|x,a))$, where each element represents the probability of delaying by exactly $d$ stages (instead of cumulative).
+For $T=4$, the delay mechanism parameters are $\bm{\rho}(\texttt{F},1)=(0.64,0.18,0.07,0.03)$ and $\bm{\rho}(\texttt{F},0)=(0.63,0.18,0.05,0.02)$ for the female subgroup, and $\bm{\rho}(\texttt{M},1)=(0.55,0.23,0.10,0.02)$ and $\bm{\rho}(\texttt{M},0)=(0.54,$ $0.11,0.21,0.01)$ for the male subgroup.
+We show in panels (A)–(C) of Figure 3 the optimized variance under the power maximization objective.
+For complete randomization, the variances of the treatment effect estimator are 7.75, 8.04, and 8.17, respectively.
+The results clearly demonstrate that the proposed fully forward-looking CARA design and its DBCD-enhanced version exhibit a smaller deviation from the oracle variance (indicated by the horizontal black dashed line); to compare, complete randomization leads to severe efficiency loss.
+Interestingly, the Neyman allocation performs reasonably well in this design, although it still leads to slightly larger variances.
+Figure 3: Optimized variance (panels A–C) and failure rate (panels D–F) for experimental horizons $T=4,5,6$. Covariate stratification is based on biological sex.
+Panels (D)–(F) in the figure collects the failure rates for both our proposed method and the ethical design.
+For complete randomization, the failure rates are 0.31,0.33, and 0.35.
+Again, our proposed method and its DBCD-enhanced version tend to perform better, and the overall failure rate there is very close to the oracle level (horizontal black dashed line).
+In contrast, failure rates tend to be considerably higher for complete randomization.
+To assess the finite-sample distributional properties of the treatment effect estimator, we present the coverage probabilities in Table 1 for the two design objectives and the three extrapolations methods for the estimated delay mechanism.
+Encouragingly, the empirical coverage is extremely close to the nominal level.
+Lastly, we demonstrate and compare the power properties of different design strategies in Figure 4: all methods control the type I error rate well when the true ATE is zero; with $\tau>0$, however, our fully forward-looking CARA design has a clear advantage as it is able to detect the non-zero treatment effect more frequently.
+In summary, this synthetic case study showcases that our proposed fully forward-looking CARA design not only provides valid statistical inference in the presence of delayed responses, but it also delivers higher estimation efficiency and enhanced failure reduction for the two design objectives, which is in stark contrast to other methods that overlook the outcome delay issue.
+Table 1: Coverage probability comparison.
 Design 1: Power Maximization
 Delay perspective Proposed CARA design DBCD + Proposed CARA design
-Conservative 0.95 (0.01) 0.94 (0.01)
-Optimistic 0.94 (0.01) 0.96 (0.01)
-Neutral 0.95 (0.01) 0.96 (0.01)
+Conservative 0.94 (0.01) 0.95 (0.01)
+Optimistic 0.95 (0.01) 0.94 (0.01)
+Neutral 0.95 (0.01) 0.95 (0.01)
 Design 2: Failure Reduction
 Delay perspective Proposed CARA design DBCD + Proposed CARA design
-Conservative 0.96 (0.01) 0.95 (0.01)
-Optimistic 0.95 (0.01) 0.94 (0.01)
-Neutral 0.94 (0.01) 0.95 (0.01)
-Table 1: Coverage probability comparison across different delay perspectives under two different design objectives.
-Figure 5 (A) – (C) shows the variance of the ATE estimator under the power maximization objective across experimental stages for different designs.
-The results show that the proposed fully forward-looking CARA design and its DBCD-enhanced counterpart exhibit a smaller deviation from the oracle variance (indicated by the dashed line) than the complete randomization design and the Neyman allocation.
-This suggests that the proposed forward-looking design achieves higher estimation efficiency.
-Furthermore, Figure 5 (A) – (C) demonstrates that the efficiency of the fully forward-looking CARA design progressively improves as the experimental stages advance.
-We conjecture that this is because, as we obtain more information about the delay mechanism, the treatment allocation approaches the optimal allocation, leading to more significant changes over time.
-In contrast, while the Neyman allocation design has a notable improvement in efficiency after Stage 1, its variance stabilizes in subsequent stages.
-This is because, after the first stage, the Neyman allocation design adjusts for the $1/2$ treatment allocation initialized in Stage 1.
-Since Neyman allocation does not account for the delay mechanism, its variance quickly stabilizes over time.
-Figure 5 (D) – (E) shows a similar trend under the failure reduction design objective.
-The fully forward-looking CARA design and its DBCD-enhanced counterpart achieve high failure reduction rates.
-Moreover, while the fully forward-looking CARA design shows a gradual decline in failure reduction as experimental stages increase, the failure reduction rates for the ethical design and the complete randomization design remain stable over time.
-This suggests that the adaptive nature of the fully forward-looking CARA design facilitates dynamically improved performance across stages.
-Lastly, our proposed fully forward-looking CARA design enhanced by DBCD overall exhibits a more robust performance.
-We conjecture that this is because the DBCD design targets the current stage treatment allocation toward the optimal treatment allocation, thereby further improving the empirical performance (i.e., estimation efficiency and failure rate reduction).
-As our design is rigorously guided by the semiparametric efficiency theory to improve the estimation efficiency of the average causal effect, we present power curves for correctly detecting the true ATEs in Figure 6.
-Figure 6 suggests that when the ATE is 0, all considered designs control the type I error rate at the nominal level.
-Across two design objectives, the proposed forward-looking design overall demonstrates higher power in detecting the treatment effect compared to other designs, further validating the estimation efficiency of our proposed design.
-To demonstrate the statistical validity of our proposed designs under different experimenters’ anticipation of delay mechanism, we present the coverage probabilities of the constructed confidence interval correctly covering the true ATE under both design objectives in Table 1.
-Table 1 shows that both the proposed fully forward-looking CARA design and its DBCD-enhanced design attain the nominal-level coverage probabilities across various delay perspectives.
-This suggests that, regardless of the perspective on delay when carrying out the experiment, the resulting ATE estimators under the proposed fully forward-looking design provide valid statistical inferences.
-In summary, this synthetic case study demonstrates that our proposed fully forward-looking designs provide valid statistical inference in the presence of delayed responses.
-Additionally, our proposed designs exhibit higher estimation efficiency under the power maximization objective and improved failure reduction rate under the failure reduction objective compared to the competing designs, further validating the theoretical results presented in Section 4.
-6 Discussion and concluding remarks
-In this manuscript, we have proposed a novel fully forward-looking CARA design to address the challenges posed by delayed outcomes.
-By accounting for response delay mechanisms to identify optimal treatment allocations, our design achieves substantial improvements in experimental efficiency and participant welfare compared to traditional CARA designs.
-Through rigorous theoretical development, we derived semiparametric efficiency bounds that account for covariate-dependent delays and demonstrated that the proposed design achieves consistent and asymptotically normal treatment effect estimators.
-Our simulation studies further validated the practical advantages of this CARA design, showcasing its ability to outperform classical methods like Neyman allocation and complete randomization in terms of statistical efficiency and ethical outcomes.
-Notably, the forward-looking design dynamically adapts to evolving data, ensuring robust performance across varying experimental stages.
-Moreover, our work provides practical tools for implementing delay-informed CARA designs, offering pathways for future studies in clinical trials and other adaptive experimental settings.
+Conservative 0.94 (0.01) 0.95 (0.01)
+Optimistic 0.96 (0.01) 0.94 (0.01)
+Neutral 0.95 (0.01) 0.94 (0.01)
+Figure 4: Power comparison for the two design objectives. Covariate stratification is based on biological sex.
+Before closing this section, we provide additional simulation evidence with a different covariate stratification, defined by biological sex and the WHO clinical stages of HIV.
+Specifically, we define four groups: $\texttt{S}_{1}$ for male subjects with WHO stages 1 and 2, $\texttt{S}_{2}$ for female subjects with stages 1 and 2, $\texttt{S}_{3}$ for male with stages 3 and 4, and $\texttt{S}_{4}$ for female with stages 3 and 4.
+For the first stratum $\texttt{S}_{1}$, the calibrated parameters are $\sigma(\texttt{S}_{1},1)=1.63$, $\sigma(\texttt{S}_{1},0)=1.87$, $\bm{\rho}(\texttt{S}_{1},1)=(0.10,0.15,0.32,0.04)$, and $\bm{\rho}(\texttt{S}_{1},0)=(0.03,0.10,0.22,0.30)$.
+For the second stratum $\texttt{S}_{2}$, the parameters are $\sigma(\texttt{S}_{2},1)=1.85$, $\sigma(\texttt{S}_{2},0)=1.90$, $\bm{\rho}(\texttt{S}_{2},1)=(0.14,0.16,0.21,0.05)$ and $\bm{\rho}(\texttt{S}_{2},0)=(0.10,0.14,0.17,0.06)$.
+For $\texttt{S}_{3}$, we adopt $\sigma(\texttt{S}_{3},1)=2.37$, $\sigma(\texttt{S}_{3},0)=1.63$, $\bm{\rho}(\texttt{S}_{3},1)=(0.10,0.19,0.21,0.14)$ and $\bm{\rho}(\texttt{S}_{3},0)=(0.11,0.16,0.19,0.09)$.
+And finally for $\texttt{S}_{4}$, we use $\sigma(\texttt{S}_{4},1)=1.73$, $\sigma(\texttt{S}_{4},0)=2.14$, $\bm{\rho}(\texttt{S}_{4},1)=(0.05,0.20,0.17,0.18)$ and $\bm{\rho}(\texttt{S}_{4},0)=(0.19,0.08,0.24,0.02)$.
+The oracle treatment allocations under the proposed design and the Neyman allocation are demonstrated earlier in Figure 2.
+We present simulation results on the optimized variances in panels (A)–(C) of Figure 5, and the failure rates in panels (D)–(F).
+Although the covariate stratification differs (cf.
+Figure 3), a consistent pattern emerges: both our proposed design and its DBCD-enhanced variant yield lower variance and enhanced failure reduction by explicitly accounting for the delay mechanism.
+To compare, the variances of the treatment effect estimator employing complete randomization are 43.96, 45.17, and 47.82, and the failure rates are 0.34, 0.32, and 0.34, respectively.
+To complement our earlier findings in Figure 4, we also report power comparisons in Figure 6.
+Figure 5: Optimized variance (panels A–C) and failure rate (panels D–F) for experimental horizons $T=4,5,6$. Covariate stratification is based on biological sex and WHO clinical stages.
+Figure 6: Power comparison for the two design objectives. Covariate stratification is based on biological sex and WHO clinical stages.
+6 Conclusion
+In this paper, we introduce fully forward-looking covariate-adjusted response-adaptive (CARA) designs that effectively address the challenge of delayed outcomes in adaptive experiments, allowing the delay mechanism to be both arm- and covariate-dependent.
+Our approach sequentially estimates the delay mechanism, which is then used to inform optimal treatment allocation.
+Through a comprehensive synthetic case study, we demonstrate that our proposed design can lead to substantial gains in both experimental efficiency and participant welfare compared to traditional CARA designs.
+We provide rigorous theoretical foundation by deriving semiparametric efficiency bounds in the presence of delayed responses and establishing the consistency and asymptotic normality of treatment effect estimators.

### v3

@@
 In clinical trials, social science field experiments, and A/B tests, participants often enroll sequentially, and responses to treatments can vary due to individual-specific characteristics.
-In these contexts, covariate-adjusted response-adaptive (CARA) designs have gained increasing popularity: by dynamically adjusting treatment assignments based on accumulated covariates and outcome information during the experiment, CARA designs can be adopted to optimize treatment allocation to improve statistical power (Hahn, Hirano and Karlan, 2011; Hu, Zhu and Hu, 2015; Blackwell, Pashley and Valentino, 2023), reduce participant exposure to less effective treatments and enhance overall welfare (Hu and Rosenberger, 2006; Hu, Zhu and Hu, 2015; Wei, Ma and Wang, 2024, 2025), and maintain statistical validity for analyzing experimental data (Zhang, Hu, Cheung and Chan, 2007; Hu and Hu, 2012; Bugni, Canay and Shaikh, 2018; Offer-Westort, Coppock and Green, 2021; Bai, Romano and Shaikh, 2022; Robertson, Lee, López-Kolkovska and Villar, 2023; Zhao, 2023; Wei, Ma and Wang, 2025; Bibaut and Kallus, 2025).
+In these contexts, covariate-adjusted response-adaptive (CARA) designs have gained increasing popularity: by dynamically adjusting treatment assignments based on accumulated covariates and outcome information during the experiment, CARA designs can be adopted to optimize treatment allocation to improve statistical power (Hahn et al., 2011; Hu et al., 2015; Blackwell et al., 2023), reduce participant exposure to less effective treatments and enhance overall welfare (Hu and Rosenberger, 2006; Hu et al., 2015; Wei et al., 2023, 2025), and maintain statistical validity for analyzing experimental data (Zhang et al., 2007; Hu and Hu, 2012; Bugni et al., 2018; Offer-Westort et al., 2021; Bai et al., 2022; Robertson et al., 2023; Zhao, 2023; Wei et al., 2025; Bibaut and Kallus, 2025).
 Nevertheless, because CARA designs adjust treatment assignments based on observed participant responses, delays in outcome observation can significantly hinder their effectiveness.
@@
 These delays often depend on both the treatment arm and participant covariates, as is common in sequentially enrolled randomized experiments.
-As an example, we revisit the study by Fahey, Njau, Katabaro, Mfaume, Ulenga, Mwenda et al.
+As an example, we revisit the study by Fahey et al.
 (2020), a randomized experiment conducted in the Shinyanga region of Tanzania between April 24 and December 14, 2018.
@@
 Figure 1: Delays in viral load blood test across treatment arms and covariate strata (biological sex).
-To our knowledge, although there has been prior work on developing valid statistical inference procedures in the presence of delayed responses in CARA (Bai, Hu and Rosenberger, 2002; Hu, Zhang, Cheung and Chan, 2008; Robertson, Lee, López-Kolkovska and Villar, 2023), no CARA designs has incorporated the delay mechanism when optimizing experimental objectives.
+To our knowledge, although there has been prior work on developing valid statistical inference procedures in the presence of delayed responses in CARA (Bai et al., 2002; Hu et al., 2008; Robertson et al., 2023), no CARA designs has incorporated the delay mechanism when optimizing experimental objectives.
 In this paper, we address the challenges that response delays pose to covariate-adjusted response-adaptive (CARA) designs.
@@
 First, we demonstrate that the classical optimal allocation strategy used in existing CARA designs no longer provides optimal treatment assignment when accounting for arm- and covariate-dependent delays.
-Specifically, we first derive a semiparametric efficiency bound for treatment effect estimation in the presence of outcome delay (Theorem 1), which differ substantially from existing results in the literature (Hahn, 1998; Hirano, Imbens and Ridder, 2003; Cattaneo, 2010; Zhu and Zhu, 2023; Wu, Zheng, Zhang, Zhang and Wang, 2025).
+Specifically, we first derive a semiparametric efficiency bound for treatment effect estimation in the presence of outcome delay (Theorem 1), which differ substantially from existing results in the literature (Hahn, 1998; Hirano et al., 2003; Cattaneo, 2010; Zhu and Zhu, 2023; Wu et al., 2025).
 See van der Vaart (2000) and Tsiatis (2006) for general discussions on semiparametric efficiency.
 This result allows us to show that the optimal delay mechanism involves a unique trajectory of treatment assignment probabilities that vary across different stages, resulting in treatment allocations with lower variance for power maximization, and enhanced failure reduction in welfare improvement designs.
-To our knowledge, this is the first design to explicitly explore delay mechanisms to optimize trial objectives, while existing literature focuses on conducting sequential tests (Hampson and Jennison, 2013), designing Bayesian adaptive designs where delay is independent of the arm and covariates (Lin, Thall and Yuan, 2020), or conducting statistical inference after experiments have concluded (Bai, Hu and Rosenberger, 2002; Hu, Zhang, Cheung and Chan, 2008).
+To our knowledge, this is the first design to explicitly explore delay mechanisms to optimize trial objectives, while existing literature focuses on conducting sequential tests (Hampson and Jennison, 2013), designing Bayesian adaptive designs where delay is independent of the arm and covariates (Lin et al., 2020), or conducting statistical inference after experiments have concluded (Bai et al., 2002; Hu et al., 2008).
 Second, since the delay mechanism is unknown a priori and remains only partially estimable during the experiment, we propose a fully forward-looking CARA design that sequentially updates treatment assignment probabilities to target a trial objective at the conclusion of the experiment.
 In this process, our algorithm offers several extrapolation strategies when estimating delay mechanism, allowing better alignment with the designer’s prior knowledge.
-We further discuss in detail two design objectives frequently adopted in current CARA literature: maximizing experiment power (Tymofyeyev, Rosenberger and Hu, 2007; Zhao, 2023) and improving overall participant welfare (or failure reduction) while maintaining statistical power constraints (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001; Robertson, Lee, López-Kolkovska and Villar, 2023).
+We further discuss in detail two design objectives frequently adopted in current CARA literature: maximizing experiment power (Tymofyeyev et al., 2007; Zhao, 2023) and improving overall participant welfare (or failure reduction) while maintaining statistical power constraints (Rosenberger et al., 2001; Robertson et al., 2023).
 We justify the benefits of our proposed forward-looking CARA design through theoretical investigation (Section 4) and a synthetic case study (Section 5).
@@
 From the above result, it is evident that when the delay mechanism of the response depends on the covariates and treatment status, the optimal treatment allocation rules in classical CARA designs must be modified to account for outcome delays.
-In what follows, we examine how the optimal treatment allocation rule is affected by the arm- and/or covariate-dependent delayed responses in two common experimental objectives: one aimed at enhancing statistical power (Tymofyeyev, Rosenberger and Hu, 2007), and the other focused on minimizing the expected number of failures while maintaining a minimum power requirement for binary outcomes (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001).
+In what follows, we examine how the optimal treatment allocation rule is affected by the arm- and/or covariate-dependent delayed responses in two common experimental objectives: one aimed at enhancing statistical power (Tymofyeyev et al., 2007), and the other focused on minimizing the expected number of failures while maintaining a minimum power requirement for binary outcomes (Rosenberger et al., 2001).
 In both design goals, the variance of the estimated treatment effects plays an essential role.
 To simplify the presentation, we focus on two-arm CARA designs with $a\in\{0,1\}$ in this paper.
-For multi-arm experiments, the design objective can be generalized using non-centrality parameters, as discussed in Lachin (1977) and Tymofyeyev, Rosenberger and Hu (2007).
+For multi-arm experiments, the design objective can be generalized using non-centrality parameters, as discussed in Lachin (1977) and Tymofyeyev et al.
+(2007).
 Design objective 1: Optimal treatment allocation for power maximization.
@@
 $\min_{e_{1}(\cdot),\dots,e_{T}(\cdot)}\ \mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))\quad\text{subject to}\quad\delta\leq e_{t}(\cdot)\leq 1-\delta\quad\text{for }t=1,\ldots,T,$ (3)
-where $\delta$ ensures that the assignment probabilities are bounded away from 0 and 1, maintaining sufficient randomness in treatment assignments (Ma and Wang, 2020; Heiler and Kazak, 2021; Sasaki and Ura, 2022; Ma, Sasaki and Wang, 2025; Dorn, 2025).
+where $\delta$ ensures that the assignment probabilities are bounded away from 0 and 1, maintaining sufficient randomness in treatment assignments (Ma and Wang, 2020; Heiler and Kazak, 2021; Sasaki and Ura, 2022; Ma et al., 2025a; Dorn, 2025).
 The objective function $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ is a component of the semiparametric efficiency bound related to the propensity scores, and is defined as:
@@
 Design objective 2: Optimal treatment allocation for failure reduction subject to a power constraint.
-This design aims to minimize the expected number of failures while satisfying a minimum power requirement for binary responses (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001).
+This design aims to minimize the expected number of failures while satisfying a minimum power requirement for binary responses (Rosenberger et al., 2001).
 In this context, a treatment is considered to fail when the potential outcome equals one.
@@
 $e_{t}^{\mathtt{Neyman}}(a|x)=\frac{\sigma(x,a)}{\sigma(x,1)+\sigma(x,0)},\quad\text{for }t=1,\ldots,T,$ (5)
-and the optimal failure reduction rule of Rosenberger, Stallard, Ivanova, Harper and Ricks (2001):
+and the failure reduction rule of Rosenberger et al. (2001):
 $e_{t}^{\mathtt{FR}}(a|x)=\frac{\sqrt{\mu(x,a)}}{\sqrt{\mu(x,1)}+\sqrt{\mu(x,0)}},\quad\text{for }t=1,\ldots,T,$ (6)
@@
 In practice, the delay mechanism is rarely known and must be estimated as data are sequentially accumulated to optimize the experimental objective.
-In this section, we propose two delay-informed practical CARA designs tailored to two previously discussed experimental goals: improving statistical power (Tymofyeyev, Rosenberger and Hu, 2007) and minimizing the expected number of failures while meeting a minimum power requirement for binary responses (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001).
+In this section, we propose two delay-informed practical CARA designs tailored to two previously discussed experimental goals: improving statistical power (Tymofyeyev et al., 2007) and minimizing the expected number of failures while meeting a minimum power requirement for binary responses (Rosenberger et al., 2001).
 We also provide tools for valid statistical inference after the experiment is finished.
@@
 Our method is designed to anticipate the impact of future allocations and outcome delays on the variance of the final treatment effect estimate after the experiment concludes:
-$\hat{\mathsf{V}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Bigg{[}\frac{\hat{\sigma}_{t}^{2}(x,1)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)\hat{e}^{*}_{\ell}(1|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)e_{\ell}(x)}$
-$\qquad\qquad+\frac{\hat{\sigma}_{t}^{2}(x,0)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)\hat{e}^{*}_{\ell}(0|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)(1-e_{\ell}(x))}\Bigg{]}.$
+$\hat{\mathsf{V}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Bigg[\frac{\hat{\sigma}_{t}^{2}(x,1)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)\hat{e}^{*}_{\ell}(1|x)+\sum\limits_{\ell=t+1}^{T}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)e_{\ell}(x)}$
+$\qquad\qquad+\frac{\hat{\sigma}_{t}^{2}(x,0)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)\hat{e}^{*}_{\ell}(0|x)+\sum\limits_{\ell=t+1}^{T}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)(1-e_{\ell}(x))}\Bigg].$
 Here, $\hat{p}_{t}(x)=N_{t}^{-1}\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,S_{i}\leq t)}$, which is the estimated proportion of participants with covariate $x$ up to stage $t$.
@@
 $\hat{\mathsf{P}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)=$ $\sum_{\ell=1}^{t}\hat{r}_{\ell}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Big[\hat{e}^{*}_{\ell}(1|x)\left(1-\hat{\mu}_{t}(x,1)\right)+\hat{e}^{*}_{\ell}(0|x)\left(1-\hat{\mu}_{t}(x,0)\right)\Big]$
-$\quad+\sum_{\ell=t+1}^{T}r_{l}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Big{[}e_{\ell}(x)\left(1-\hat{\mu}_{t}(x,1)\right)+(1-e_{\ell}(x))\left(1-\hat{\mu}_{t}(x,0)\right)\Big{]}.$
+$\quad+\sum_{\ell=t+1}^{T}\hat{r}_{\ell}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Big[e_{\ell}(x)\left(1-\hat{\mu}_{t}(x,1)\right)+(1-e_{\ell}(x))\left(1-\hat{\mu}_{t}(x,0)\right)\Big].$
 We denote the solution to the following optimization problem as $(\tilde{e}_{t+1}(\cdot),\dots,\tilde{e}_{T}(\cdot))$:
@@
    Report the $(1-\alpha)\%$ confidence interval: $\hat{\tau}_{T}\pm z_{1-\frac{\alpha}{2}}\sqrt{\hat{\mathsf{V}}_{T}/N}$.
-In the following sections, we will provide theoretical guarantees and simulation evidence to demonstrate the advantages of our proposed fully forward-looking CARA designs in the presence of delayed outcomes.
 4 Theoretical Investigation
-This section provides theoretical justifications for the proposed forward-looking CARA design.
+This section provides theoretical justifications for the proposed CARA design.
 To begin, we present the semiparametric efficiency bound for treatment effect estimation in a multi-stage setting with delayed outcomes.
@@
 $\mathsf{V}=\sum_{x\in\mathcal{X}}p(x)\Bigg[\frac{\sigma^{2}(x,1)}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|x,1)e_{t}(1|x)}$ $+\frac{\sigma^{2}(x,0)}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|x,0)e_{t}(0|x)}+\big(\tau(x)-\tau\big)^{2}\Bigg].$
-We note that while the previous theorem is developed in the context of discrete covariates (with the efficiency bound expressed as a summation over the probability mass function), it can be generalized to accommodate, for example, continuously distributed covariates (c.f., equation 1).
+We note that while the previous theorem is developed in the context of discrete covariates, it can be generalized to accommodate continuously distributed covariates (c.f., equation 1).
 4.2 Asymptotic Properties of Treatment Effects Estimates
 In this subsection, we demonstrate that the treatment effect estimator is consistent and admits an asymptotic normal distribution.
-The main conclusions of this section rely on a high-level condition regarding the convergence of the optimized treatment allocation rule, making these conclusions applicable to a wide range of CARA and a broad class of frequentist adaptive experimental design settings (Hu and Zhang, 2004; Antognini and Zagoraiou, 2015; Hu, Zhu and Hu, 2015).
+The main conclusions of this section rely on a high-level condition regarding the convergence of the optimized treatment allocation rule, making these conclusions applicable to a wide range of CARA and a broad class of frequentist adaptive experimental design settings (Hu and Zhang, 2004; Antognini and Zagoraiou, 2015; Hu et al., 2015).
 We begin with this high-level condition, which will be verified in the next subsection for our proposed forward-looking CARA design.
@@
 That said, it is worth clarifying that, in general, one cannot achieve the optimized oracle efficiency bound, an issue we discuss further below.
-To conclude this subsection, the following result establishes valid statistical inference, where the variance estimator $\hat{\mathsf{V}}_{T}$ is defined at the end of Section 3.
 Theorem 3 (Statistical inference).
@@
 For this reason, we have introduced feasible and concrete ways to extrapolate the delay mechanism.
-Following our discussion in Section 2, we define
-$\rho^{\dagger}_{t}(d|x,a)$ $=\min\big{\{}\rho(d|x,a),\rho(t-1|x,a)\big{\}}.$
+Following our discussion in Section 2, we define $\rho^{\dagger}_{t}(d|x,a)=\min\big\{\rho(d|x,a),\rho(t-1|x,a)\big\}.$
 It is helpful to compare this definition with Assumption 2.
@@
 5 Synthetic Case Study
-To evaluate the performance of the proposed forward-looking CARA design, we conduct simulation studies using a data generation mechanism calibrated from the study of Fahey, Njau, Katabaro, Mfaume, Ulenga, Mwenda et al.
+To evaluate the performance of the proposed forward-looking CARA design, we conduct simulation studies using a data generation mechanism calibrated from the study of Fahey et al.
 (2020).
-Their original experiment aims to evaluate the effectiveness of cash transfers on retention in care among people living with HIV (PLHIV), using viral load as the primary response variable.
-The experiment participants were adults aged 18 years or older diagnosed with HIV, who were randomized to receive cash transfers (treatment arm, $A_{i}=1$) or no cash transfer (control arm, $A_{i}=0$).
-The original study involved two treatment arms with different levels of cash incentives, which are combined into one group in our synthetic case study.
-The response variable, viral load, was assessed six months after enrollment.
-Although the original study was not conducted in an adaptive fashion, adaptation can be applied since the participants were enrolled sequentially.
-However, a challenge is that the primary responses of the participants were collected with delays, with some responses collected longer than six months after the treatment assignment; see Figure 1.
-Our synthetic case study conducts a simulation experiment in which the underlying data-generating process (delay mechanism, covariate distribution, and potential outcome distributions) is calibrated with estimates from the original study.
-For the first set of simulation results, we consider biological sex as the only covariate $X_{i}$, leading to two subgroups: $X_{i}=\mathtt{M}$ and $X_{i}=\mathtt{F}$ for male and female participants.
-From the original data, the distribution of biological sex consists of 36% male and 64% female, which we adopt as the data-generating process of $X_{i}$ in our Monte Carlo exercise.
+For the first set of simulation results, we consider biological sex as the only covariate $X_{i}$, leading to two subgroups: $X_{i}=\mathtt{M}$ and $X_{i}=\mathtt{F}$ with frequencies 36% and 64%, respectively.
 We set $n_{t}=100$ for $t=1,\ldots,T$ with $T\in\{4,6,8\}$.
@@
 Setup 1: Power maximization.
-For the power maximization design objective, we compare four design methods: (i) our proposed design; (ii) our design enhanced by the doubly adaptive biased coin design (DBCD) (Hu and Zhang, 2004; Tymofyeyev, Rosenberger and Hu, 2007); (iii) complete randomization; (iv) Neyman allocation.
-We note that DBCD is a response-adaptive randomization design that more robustly adjusts treatment allocation toward the optimal allocation dynamically.
+For the power maximization design objective, we compare four design methods: (i) our proposed design; (ii) our design enhanced by the doubly adaptive biased coin design (DBCD) (Hu and Zhang, 2004; Tymofyeyev et al., 2007); (iii) complete randomization; (iv) Neyman allocation.
+We note that DBCD is a response-adaptive randomization design that adjusts treatment allocation toward the optimal allocation dynamically.
 Since our approach involves assigning treatments across multiple covariate strata, we apply the DBCD design separately within each stratum.
@@
 Setup 2: Failure reduction.
-For the failure reduction objective, we again compare across four designs: (i) the proposed method; (ii) our proposed design again enhanced by DBCD; (iii) the ethical design proposed by Rosenberger, Stallard, Ivanova, Harper and Ricks (2001); (iv) complete randomization.
+For the failure reduction objective, we again compare across four designs: (i–ii) the proposed method (with and without DBCD enhancement); (iii) the ethical design (Rosenberger et al., 2001); (iv) complete randomization.
 In this simulation setup, we transform the original viral load measurements into a binary variable, indicating whether the participant has achieved viral suppression; specifically, if a participant’s viral load is below $1,000$ copies per mL, we set $Y_{i}=1$ (viral suppression).
@@
 Encouragingly, the empirical coverage is extremely close to the nominal level.
-Lastly, we demonstrate and compare the power properties of different design strategies in Figure 4: all methods control the type I error rate well when the true ATE is zero; with $\tau>0$, however, our fully forward-looking CARA design has a clear advantage as it is able to detect the non-zero treatment effect more frequently.
-In summary, this synthetic case study showcases that our proposed fully forward-looking CARA design not only provides valid statistical inference in the presence of delayed responses, but it also delivers higher estimation efficiency and enhanced failure reduction for the two design objectives, which is in stark contrast to other methods that overlook the outcome delay issue.
+Lastly, we demonstrate and compare the power properties of different design strategies in Figure 4: all methods control the type I error rate well when the true ATE is zero; with $\tau>0$, however, our CARA design has a clear advantage as it is able to detect the non-zero treatment effect more frequently.
+In summary, this synthetic case study showcases that our fully forward-looking CARA design not only provides valid statistical inference in the presence of delayed responses, but it also delivers higher estimation efficiency and enhanced failure reduction for the two design objectives.
 Table 1: Coverage probability comparison.
@@
 Figure 4: Power comparison for the two design objectives. Covariate stratification is based on biological sex.
-Before closing this section, we provide additional simulation evidence with a different covariate stratification, defined by biological sex and the WHO clinical stages of HIV.
+Before closing this section, we provide additional simulation evidence where stratification is based on biological sex and the WHO clinical stages of HIV.
 Specifically, we define four groups: $\texttt{S}_{1}$ for male subjects with WHO stages 1 and 2, $\texttt{S}_{2}$ for female subjects with stages 1 and 2, $\texttt{S}_{3}$ for male with stages 3 and 4, and $\texttt{S}_{4}$ for female with stages 3 and 4.
@@
 The oracle treatment allocations under the proposed design and the Neyman allocation are demonstrated earlier in Figure 2.
-We present simulation results on the optimized variances in panels (A)–(C) of Figure 5, and the failure rates in panels (D)–(F).
+We present simulation results on the optimized variances in panels (A)–(C) of Figure 5, and the failure rates in (D)–(F).
 Although the covariate stratification differs (cf.
-Figure 3), a consistent pattern emerges: both our proposed design and its DBCD-enhanced variant yield lower variance and enhanced failure reduction by explicitly accounting for the delay mechanism.
+Figure 3), a consistent pattern emerges: our proposed design yields lower variance and enhanced failure reduction by explicitly accounting for the delay mechanism.
 To compare, the variances of the treatment effect estimator employing complete randomization are 43.96, 45.17, and 47.82, and the failure rates are 0.34, 0.32, and 0.34, respectively.
@@
 We provide rigorous theoretical foundation by deriving semiparametric efficiency bounds in the presence of delayed responses and establishing the consistency and asymptotic normality of treatment effect estimators.
+Supplementary Material
+The Supplementary Material (Ma et al., 2025b) contains general theoretical results, their proofs, and reports additional simulation evidence.
+Acknowledgments
+We thank the editor, the associate editor, two anonymous reviewers, seminar participants, and conference audiences for their valuable feedback, which significantly improved the manuscript.
+Xinwei Ma gratefully acknowledges financial support from the National Institutes of Health through the grant R01-AG089512.
+Jingshen Wang gratefully acknowledges financial support from the National Institutes of Health through the grant R01-AG089512, and the National Science Foundation through the grants DMS-2239047 and DMS-2220537.